### v2

# Superconducting qubits at the utility scale: the potential and limitations of modularity

Abstract
The development of fault-tolerant quantum computers (FTQCs) is receiving increasing attention within the quantum computing community. Like conventional digital computers, FTQCs, which utilize error correction and millions of physical qubits, have the potential to address some of humanity’s grand challenges. However, accurate estimates of the tangible scale of future FTQCs, based on transparent assumptions, are uncommon. How many physical qubits are necessary to solve a practical problem intractable for classical hardware? What costs arise from distributing quantum computation across multiple machines? This paper presents an architectural model of a potential FTQC based on superconducting qubits, divided into discrete modules and interconnected via coherent links. We employ a resource estimation framework and software tool to assess the physical resources required to execute specific quantum algorithms compiled into their graph-state form and arranged onto a modular superconducting hardware architecture. Our tool can predict the size, power consumption, and execution time of these algorithms as they approach the utility scale based on explicit assumptions about the physical layout, thermal load, and modular connectivity of the system. Using this tool, we assess the total resources in the proposed modular architecture and highlight the impact of trade-offs, intermodule connectivity, latency, and space-time resource requirements.
Rapid advancement in the development of FTQCs heralds a new era in which universal machines with millions of qubits can surpass classical computers in addressing specific problems (see [1, 2, 3, 4, 5, 6] for recent overviews and developments). At the same time, the size, connectivity, and fidelity of today’s noisy qubit devices are continuously improving [7, 8, 9]. These platforms are already considered the foundational elements of early FTQCs. They employ redundancy to suppress noise using a technique known as quantum error correction (QEC) [10, 11, 12, 4, 13, 9].
The complexity and asymptotic scaling of specific fault-tolerant algorithms, including Shor’s factoring [2, 6], quantum signal processing [14] (QSP), and ground state energy estimation [15] (GSEE), are well understood. However, research on the magnitude and scaling of the physical resources needed to execute such computations is limited. Furthermore, the overhead and space-time resource requirements for physical hardware are even less well understood and depend on various choices related to platforms and micro-architectures. Scaling tangible resources such as energy, time, and space based on specific assumptions about a potential machine is crucial for evaluating the feasibility of methods to develop FTQC.
Here, we present a transparent fault-tolerant architectural model based on square lattices of superconducting qubits utilizing the surface code [10, 12, 16, 17, 18], allocated to the essential elements of FTQC: logical Clifford gates through lattice surgery, magic or $T$ state distillation (MSD) [11, 12, 19] for non-Clifford gates, as well as considering the queuing and routing of specific resources. We also establish explicit assumptions regarding the size of a single module and the penalties associated with executing a fault-tolerant algorithm that exceeds the capacity of a single module to quantify the impact of coherently distributing a quantum computation.
Resource estimation tools exist based on the “T-counting” method [12, 20, 21, 22, 23]. These footprint-style estimation tools synthesize the logical circuit into Clifford+$T$ gates (or other universal gate sets), rescale the space-time volume according to the desired level of parallelization, and derive physical space-time quantities based on the $T$-count and circuit depth, along with assumptions regarding $T$-distillation and hardware timescales. This methodology lacks targeted estimates specific to Quantum Processing Unit (QPU) layouts, provides insufficient estimates for surface code distance, and may overestimate the upper limits of space-time resources.
We also present application-specific resource estimates for the proposed fault-tolerant superconducting qubit architecture. These estimates are based on a graph state method detailed in section I.1. This framework aims to assist in designing and assessing the feasibility of next-generation FTQCs while quantifying the trade-offs associated with hardware implementations. We also discuss an accompanying open-source software called Rigetti Resource Estimations [24] (RRE), which we used to generate all the results presented in section III for a selection of test algorithms outlined in section II. We will explore potential future work in section IV and provide a conclusion in section V.
I A primer for algorithm execution, architecture designs, and resource estimation on FTQCs
The execution of algorithms on an FTQC fundamentally differs from merely executing the gates of a quantum circuit in the order they appear. By definition, FTQCs depend on the underlying QEC [10, 11, 12, 4, 13, 9] scheme they utilize. FTQCs employ noisy physical qubits, encoded within an error correction scheme, to create logical qubits with exponentially suppressed error rates.
The error-correcting scheme we consider here is the surface code [10, 12, 16, 17], which is implemented on square arrays of physical qubits. In an array, qubits are categorized into physical data or ancilla qubits, with each connected data qubit coupled to four ancilla qubits and vice versa. Ancilla qubits facilitate repeated composite parity measurements on their four neighboring data qubits of form $X_{1}X_{2}X_{3}X_{4}$ or $Z_{1}Z_{2}Z_{3}Z_{4}$. The composite parity measurements on the data qubits can be performed by entangling them with the ancilla through CNOTs and measuring the ancilla in either the Pauli $X$ or $Z$ basis. Physical errors on qubits in these arrays result in patterns of bit flips among the parity measurements known as syndromes. Tracking the outcomes of the parity measurements, the resulting error syndromes, and mapping them to the most probable error event is known as decoding. In practical terms, the corrections applied after decoding can be accomplished through judicious syndrome tracking and implementing corrections in software after running the algorithm (potentially incurring time delays due to decoder latency).
The price we pay for this encoding is that the resulting logical qubits can only perform a limited set of operations, fault-tolerantly, specifically the Clifford set. At least one non-Clifford operation must be engineered in the FTQC to maintain universality and execute arbitrary fault-tolerant algorithms. Typically, and for this paper, we consider the $T$ gate: a $\pi/4~{}R_{z}$-rotation. However, forcing an error-correcting code to produce an operation outside its native subspace, like a $T$-gate, involves significant overhead, necessitating the dedication of substantial portions of the FTQC to MSD or, more recently, cultivation [11, 12, 19].
Figure 1: (top) Fault-tolerant algorithm execution involves segmenting the application into logical Clifford and non-Clifford portions. Clifford portions are accomplished through a set of measurements of logical qubits facilitated by an ancilla bus. Due to the probabilistic nature of measurement outcomes, classical feedback and feedforward are incorporated between the schedules of these operations. Non-Clifford portions are accomplished using magic state distillation. (bottom) This results in a layout of the logical micro-architecture shown for illustration purposes only. Each square represents a surface code patch of distance $d$. A register of logical memory qubits (blue tiles) connects through a bus of ancilla qubits (yellow tiles). Below it, a separate bus is designated for queuing $T$-states created in an array of $T$-factories. This architecture can be distributed across multiple machines, provided they are connected via the ancilla bus and each module consumes local $T$-states. We implement a modified version with a wrapped, more practical quantum register, as seen in fig. 6.
Due to the restrictions imposed by the error-correcting code, we ultimately divide the algorithm into operations consisting of Clifford+$T$ gates executed sequentially. This separation of functional operations leads to a specialization of the fundamental architecture: Certain sections of the FTQC are dedicated solely to performing the original logical Clifford operations through multi-product (qubit) Pauli measurements—a process called lattice surgery [25, 26, 27, 28, 12]. For example, the multi-product Pauli measurements can be understood as executing $n$-body $CZ$ and single-qubit Clifford gates on the quantum state, leading to a formalism we later explore. Additionally, there are designated areas for generating and queuing $T$ states produced via MSD. An example of this layout is shown schematically in fig. 1.
We adapt this logical quantum operation model to a fixed yet extensible superconducting qubit architecture, modified from fig. 1 and consisting of a physical layout and square-lattice connectivity. We estimate the execution of selected algorithms on this architecture. To facilitate this, we generate schedules of quantum operations using a graph-state-based approach, as proposed in several studies [29, 30, 31, 32, 33, 34, 35, 36, 37] and further refined below. This approach yields several abstractions that aid in the distribution of the algorithm across the fixed hardware. Namely, we can manipulate a graph or subgraphs to customize the entire algorithm execution on an FTQC composed of distinct yet interconnected modules.
One can implement the original quantum algorithm using the algorithm-specific graph (ASG) formalism [38, 31, 32, 39] and associated schedules of operations [34, 33], which apply only logical measurements and classical feedforward. The ASG formalism is a cost-efficient variant of the Measurement-Based Quantum Computing (MBQC) framework — the latter was introduced in [35, 40]. MBQC is particularly appealing for distributed FTQC because it requires only an algorithm-specific resource, such as the graph states and measurements on logical qubits. We will explore and utilize a specific form of MBQC as outlined in [31] for compilation purposes. We prepare the graph-state resource by using lattice surgery to perform entangling operations between qubits on a register of memory qubits via a single bus of ancillary qubits. Due to the probabilistic nature of quantum measurements, we must track the set of measurement results classically using a process called Pauli frame tracking [41, 42, 43, 33]. Next, we teleport local $T$-states to the register of memory qubits to apply the non-Clifford portion in conjunction with the application of the Pauli Frame corrections of the algorithm. Finally, we apply any corrections flagged by syndromes tracked by the decoder, resulting from physical errors in the hardware. We present a mathematical description of the ASG formalism in section I.1.
How can we arrange the elements of the graph-state-based execution of the algorithm across physical components at scale? In large-scale classical computations, there are limits to the amount of computation that can be executed on a single processor within a specific time frame. These limits arise from the processing power that can be integrated into an individual chip, ultimately influenced by physical constraints and manufacturing yields. Consequently, computations are often distributed across multiple processors and machines. This distribution incurs penalties, such as higher latency or reduced bandwidth, but it enables us to leverage increased overall computational power. We can draw an analogy to distributed FTQC, which involves a network of interconnected compute modules. Manufacturing yields limit the quantum computing power of each module. These operational (thermal) and manufacturing constraints restrict the number of qubits that can be incorporated into a single QPU housed within a refrigeration module. This prompted us to propose a logical macro-architecture composed of several modules described in section I.2.
In contrast to the flexible connectivity and relatively slow speeds found between compute modules, QPUs offer fixed native connectivity and faster gate speeds. In section I.3, we describe hardware micro-architecture, how it can be customized to accommodate specific hardware constraints, and a physical rationale for why the number of qubits in a single compute module should be around one million. We propose sparsely interconnecting the modules, which may result in reduced fidelity, lower bandwidth, or slower entangling gate speeds. We will refer to the connection between modules with the surface code distance of $d$ as a single "pipe." We will abstract the transduction efficiency as an interaction time, characterized by the "intermodule tock," and quantify how this time impacts the algorithm’s overall runtime (we assume that the pipes share identical noise characteristics and code distance with the rest of FTQC).
Finally, we present our compilation and estimation methodology based on user inputs in RRE in section I.4.6. Further details regarding the output glossary of RRE, our circuit decomposition strategy, and the verification of RRE outputs can be found separately in section A.1, section A.2.4, and section A.3, respectively.
I.1 Mathematical description of the graph-state formalism
A mathematical graph, denoted as $\mathcal{G}$, can be defined by its vertices and edges. The graph state associated with a quantum (Clifford) state is constructed by assigning each vertex to the qubits initialized to $\ket{+}$ and each edge to an entangling $CZ$ operation that links the qubits on the corresponding vertices. We often refer to the qubit corresponding to a vertex as a "node" and interchange the terms "graph" and "graph state" for brevity. The nodes will undergo non-Pauli basis measurements corresponding to the original circuit’s non-Clifford gates and some $X$-basis measurements to facilitate the teleportation of unknown inputs into the graph state [31].
The ASG formalism provides measurement schedules for two main stages to fully execute the fault-tolerant algorithm: one schedule details the preparation of the graph through lattice surgery (corresponding to the original nontrivial Cliffords, first stage, in fig. 1(top)) and another that outlines the necessary single-qubit measurements in non-Clifford bases (corresponding to the original non-Cliffords, second stage, in fig. 1(top)). As discussed previously, we also need to track the Pauli frame corrections in the software. We denote the nested set of stabilizer measurement schedules that "prepare" the graph in the first stage as $S^{\text{prep}}_{\mathcal{G}}$. Conversely, we denote the nested set of non-Clifford measurement schedules that "consume" the graph state in the second stage as $S^{\text{consump}}_{\mathcal{G}}$. Both schedules consist of subsets (subschedules) that specify the order of measurements, with each subset containing the nodes designated for simultaneous measurements [44, 45, 31]. Furthermore, the essential stages of MSD, routing $T$-states, and decoding can occur or be postponed at either of the two stages, depending on the requirements.
At a high level, the ASG formalism requires ingesting the fault-tolerant algorithm as single- and two-qubit gates. We employ a straightforward method for decomposition, compilation, and execution of the algorithm, segmenting it into equal-width, recurring subcircuits along the time axis. These subcircuits are referred to as "widgets," have their own (sub)graphs, and reconstruct the original algorithm when "stitched" together (see also section I.4.6). We refer to this decomposition process as "widgetization," detailed in section A.2.4. We will examine the graph properties of these widgets and the scheduling process for preparing and consuming them on the proposed hardware architecture.
I.2 High-level macro-architecture for modular graph state processing
This section describes a macro-architectural model utilizing superconducting qubits, enabling us to estimate the resources necessary for executing the fault-tolerant algorithm while understanding the associated trade-offs.
We consider a physical layout for FTQC consisting of sets of rotated surface code patches, which are tiled together to create all intra-module components (e.g., a group of patches becomes a graph-state memory register, and others become an MSD factory). Each patch contains $2d^{2}$ physical qubits. While the rotated surface code only requires $12d^{2}-$ 1 physical qubits [12, 16, 17, 18], we adapt a physical size of $2d^{2}$ for all patches to streamline our scaling calculations.
Executing the quantum algorithm based on the ASG formalism and time-direction widgetization inherently results in a macro-architecture that divides the hardware into sets of modules. The question is how to lay out modules and assign them to potentially interleaving operations. During a steady state, the execution of the quantum algorithm comprises three main steps: 1) preparation or creation of the graph, which initializes a quantum state based on the original entangling operations, 2) consumption of the graph, which executes the original non-Cliffords, and, 3) following every consumption step, the handover of the graph through Bell-pair teleportations or stitching of output nodes from the preceding widget to the input nodes of the current widget.
Figure 2: The high-level graph state processing cycle: the algorithm is divided into time-sliced widgets interleaved across two module sets in the time direction. Modules are arranged over two legs, A and B, of a ladder structure (in this example, each leg of the ladder contains three modules). Each module set manages graph preparation, consumption, and the local magic state distillation. Light purples mean the module distills, routes, or consumes $T$-states. The algorithm interconnects module sets A and B through the teleportation of output nodes from previous widgets in one leg to the input nodes of the next ones in the other leg.
Figure 3: The high-level macro-architecture consists of two sets of modules arranged like the legs of a ladder structure. Following the example in fig. 2, each leg of the ladder contains three modules. Portions of the algorithm are segmented and executed solely on individual modules along the same leg, where intra-modular interactions, characterized by high connectivity and fast gate times, dominate computations. All modules communicate through coherent links at a reduced rate and with lower connectivity, exclusively for graph preparation and handover.
Step 2 for the current widget and Step 1 for the next can co-occur in the graph state processing. This interleaved execution structure enables us to explore a degree of parallelism in a distributed FTQC. We propose distributing the interleaving preparation and consumption steps into two physically distinct sets of modules positioned along the two legs of a ladder structure. The graph processing cycle is streamlined because each set of modules alternates between only preparing or consuming the graph at each step. In contrast, the other sets of modules do the opposite on the adjacent widget, as illustrated in fig. 2. During all preparation and consumption steps, the FTQC performs MSD within the modules and fills in a dedicated $T$-transfer bus as required. The timescales for $T$-state distillation and inter-module operations are typically one order of magnitude worse than other surface code operations, including graph preparations, so MSD will remain local to improve the speed of computations. Only the consumption stage requires the FTQC to apply the $T$ states; since the graph preparations and MSD are performed independently, we need to account for possible preparation, MSD, and other delays in the consumption stage of each widget. For the consumption steps, some graph-state nodes need a single $T$-basis measurement, while others require multiple $T$-basis measurements to execute non-Clifford rotations in the original algorithm. Subsequently, FTQC teleports the current graph’s output nodes to the newly prepared widget’s input nodes. This cycle continues until FTQC processes all widgets in the complete algorithm (see also section I.4.6).
The architecture must be scalable for algorithms requiring more space-time resources than a single module or time step. Furthermore, MSD is arguably the most resource-intensive component of a universal FTQC, and we set an equal number of MSD factories supplying the $T$ states locally in each module. To address both issues, we propose a high-level ladder-style macro-architecture illustrated in fig. 3 (see fig. 6 for the intra-module layout). The quantum memory register will be shared equally among all modules on the same leg of the ladder at any given time. Here, we can utilize the faster native connectivity of a lattice of superconducting qubits within a module for graph consumption and MSD, while reduced connectivity or slower interactions are employed for some graph preparation operations and teleportation of the output nodes of graph states on one leg of the ladder to the input nodes on the other leg.
We can estimate the time required to execute the complete schedule of operations across the system by considering all local and inter-module operations. To account for the inter-module operations, we abstract away the characteristics of the teleportation interaction between modules as a single time referred to as intermodule handover time and set it to $t_{\text{inter}}=1~{}\mu$s. We incorporate this characteristic time to measure the cost of scheduled operations crossing a module boundary, estimating the total fault-tolerant handover time.
I.3 Superconducting hardware micro-architecture and lower-level components
We propose a micro-architecture concept (see fig. 6 for an illustrative example) that informs a system-level model, which must, at a minimum, include the following components:
  1. 1.
    A lattice of rotated surface code qubits used as logical patches for building all components,
  2. 2.
    A register of logical graph-state or memory qubits for graph preparation and consumption,
  3. 3.
    Continuous buses of logical ancilla qubits that couples with, connects to, and operate on memory qubits while routing $T$-states for graph consumption.
  4. 4.
    A continuous bus of ancillary qubits for queuing $T$-states must maintain a broad enough connection to the preceding ancillary qubits to teleport $T$-states to them on demand.
  5. 5.
    $T$ (or MSD) factories.
  6. 6.
    Quantum interconnects between modules to enable connections along the ancillary bus portions.
Section I highlighted the critical mechanism that executes quantum algorithms in ASG formalism: a series of non-Clifford rotations and multi-product Pauli measurements on logical graph-state qubits, supported by the ancillary buses, perform logical operations. Consequently, this architecture requires cryostat modules to connect solely along the buses of ancilla qubits. This significantly simplifies the hardware, as dense, long-range connectivity between modules is unnecessary. The essential requirement is that the number of coherent links along the buses must equal the code distance, $d$; adding additional connections will enhance overall computation for specific applications. We will quantify this trade-off in section III.
We must outline the system’s physical specifications to account for the necessary resources to execute the algorithm. We start with the layout of the qubits. To provide the most relevant estimates for current hardware, we base our projections on a superconducting qubit architecture that consists of tunable transmon qubits coupled through tunable couplers [46]. This architecture has been implemented with high fidelity by several groups [47, 48, 49, 50]. Here, we lay out the transmon qubits in a square lattice coupled to one another via tunable couplers. This square lattice directly corresponds to the surface code, allowing us to relate qubit numbers to the code distance easily.
Figure 4: Qubits coupled through tunable couplers in a surface code layout. Each qubit and coupler requires a single control line along with their respective ground connections. This limits the pitch of the qubits to match that of the signal lines, fixing the qubit-to-qubit pitch at four times the signal line pitch.
To determine the number of qubits integrated into a single module and the overall size of the QPU, we must assess the pitch of the qubits. In this tunable coupled transmon architecture, each qubit and coupler requires at least one control line[51] in addition to an extra microwave line shared among qubits for multiplexed readout[52]. While most QPUs employ perimeter wire bonding for individual control lines, several proposals and demonstrations focus on addressing qubits individually using through-substrate vias (TSVs)[53, 54].
By adopting the TSVs method, QPUs require one via per qubit and one per tunable coupler. Each signal line needs at least one ground line for return currents. As schematically illustrated in fig. 4, achieving a sufficient number of signal and ground connections is possible if the qubit pitch $P_{q}$ is four times the via pitch $P_{v}$. Ultimately, the transmon, coupler, and TSVs can all be fabricated with a pitch of less than $100~{}\mu\text{m}$. However, the bottleneck occurs at the input and output: each signal must interface with wiring and, eventually, control electronics. Any significant fan-out near the qubit chip would be costly; therefore, we assume a compliant joint to establish a reusable interconnect with a printed circuit board. With this assumption, a $P_{v}=250~{}\mu\text{m}$ interposer would result in a $P_{q}=1~{}\text{mm}$. This leads to a qubit density of one million qubits/m2 assumed in our architectural model.
Fabricating a single device that contains hundreds of thousands, or even millions, of physical qubits may seem unrealistic by today’s standards, but we can assess where the fundamental limitations lie. Most features in a superconducting qubit circuit extend beyond the micron scale, where existing processes for flat panel displays can easily handle critical dimensions. Consequently, the distributed and lumped-element microwave circuitry and TSVs could likely be produced on substrates at the m2 scale. However, the most fundamental component of the transmon qubit is the Josephson junction (JJ) [55], which is often defined using electron beam lithography [56]. In this context, yield and manufacturing precision are crucial for frequency targeting and other factors. While recent advancements have been made in tuning the properties of JJs, which have led to more scalable manufacturing methods [57, 58], adapting an appropriate JJ fabrication process at the m2 scale may prove unattainable.
Figure 5: Qubit chips are tiled on a carrier. Tunable couplers manage two-qubit gates on qubits within the chip. ICC refers to "inter-chip coupler".
In contrast, tiling superconducting dies on a single monolithic carrier is feasible, necessitating the execution of high-fidelity two-qubit gates across the interface between adjacent qubit dies. This concept has been demonstrated in [8], where a functionally similar tunable coupler [59], commonly used in high-performing transmon qubit systems, can bridge the gap between chips. Based on this, we develop our architectural model on qubit chips tiled on a common carrier, featuring native four-fold connectivity across the entire carrier fig. 5. While other yield concerns exist regarding the application of multi-chip modules, as noted in [60], we assume dies to be pre-screened and defect-free for this paper.
I.3.1 Decoder and noise modeling
To conduct physically grounded, device-specific resource estimations, we have developed an error-scaling pipeline to integrate the effects of decoding and the noise properties of superconducting systems. This pipeline converts a device-level noise model into the physical-to-logical error scaling law used in our resource counts. As part of our pipeline, we selected the square planar iSWAP surface code [16] as implemented by the midout package [61]. To generate input noise models for the QEC simulations, we create physical-device-inspired two-qubit correlated noise models for our $iSWAP$ gates. We then apply the resultant terms in a Pauli error model to simulate the noise process in a standard Clifford Tableau simulator, Stim [62]. For simplicity in this initial workflow, we use a Minimum Weight Perfect Matching (MWPM) decoder implemented in pymatching [63].
We assume a power-law type physical error to logical error scaling law for a single cycle of the surface code [10, 64]:
$p_{C}=\kappa\left(\frac{p}{p_{\text{thresh}}}\right)^{(d+1)/2},$ (1)
where $p_{C}$ is the per-cycle logical error rate and $p$ denotes the physical error rate. $\kappa$ and $p_{\text{thresh}}$ become fitting parameters. As part of our MWPM decoder simulations and device noise modeling, we sweep through the $d$-values consuming one and two qubit error channels, which led to $p_{C}[\text{MWPM}]\approx 0.52\left(\frac{p}{0.14}\right)^{\frac{d+1}{2}}$, i.e., $\kappa[\text{MWPM}]=0.52$ and $p_{\text{thresh}}[\text{MWPM}]=0.14$. The scaling coefficient values serve as examples and reflect our choice of simulation configuration. The resource estimates that follow are commensurate with this choice. RRE allows the user to specify the coefficients appropriate to their device - see section IV for an alternative approach.
I.3.2 Intra-module architecture
We aim to estimate the resources necessary to execute the algorithm on the FTQC depicted in fig. 3, comprising two sets of modules. This section presents the intra-module architecture, which is identical for all square-grid modules and used for all resource estimates. The architecture is illustrated in fig. 6 (modified from fig. 1).
Figure 6: The proposed intra-module architecture, which includes graph preparation, consumption, local magic state distillation, $T$, and Bell states queuing with some parallelization. We arrange logical qubits on a square lattice of surface code patches made of physical qubits. A left portion of the module is allocated to support logical memory and ancillary qubits for graph preparation, consumption, $T$, and Bell states routing. The continuous $T$-transfer bus is positioned in the middle to queue and teleport $T$-states distilled by magic state distilleries on the opposite side of the module.
Each module must manage graph state preparation, consumption, local MSD, $T$, and Bell states routing. Within each module, the FTQC executes the algorithm based on the ASG formalism through intra- and inter-module operations on a wrapped logical memory register (blue tiles), mediated by continuous combs of logical ancillary qubits (yellow and light yellow tiles in fig. 6). The blue quantum memory chain is part of the complete memory register shared among all $n^{\text{per-leg}}_{\text{modules}}$ modules on the same leg of the macro-architecture. At any moment, the ancillary bus portions on the same leg can communicate through $d$-wide interconnects. The yellow ancilla bus mediates logical measurements from the schedules on most of the blue memory qubits. In contrast, the light yellow supplementary bus performs the same task on the otherwise inaccessible corner memory tiles and routes Bell pairs required for the graph handover. We assign two connected ancilla and supplementary buses to simplify the resource scaling calculations, maintain the symmetry of intra-module patterns, and highlight their roles; they are otherwise identical ancillary bus elements. More efficient configurations may exist for embedding the intra-module components depending on the problem requirements–fig. 6 was selected solely for its simplicity and scalability.
The quantum memory register over all modules on the same leg has a logical size of $\tilde{n}_{\text{logical}}$ (totaling $2\tilde{n}_{\text{logical}}$ patches for the macro-architecture). Modules have a designated area comprised of a minimum $2\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$ patches for the blue memory register and yellow ancilla bus portions. We allocate one complete column and one connected tile to the supplementary bus ($l_{\text{edge}}+1$ patches on the far left side in fig. 6), then assign the memory register and ancilla bus portions to one complete edge of an individual module. The memory register and ancilla bus portions form a symmetric bilinear comb pattern repeated throughout the modules. The bilinear pattern occupies as much space as necessary to support the graph state processing on $\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$ memory nodes following the schedules. The last module structures, which may require fewer active patches for the memory register and ancillary bus portions, remain unchanged for consistency.
The FTQC performs multi-product Pauli measurements of $\tilde{S}^{\mathcal{G}}_{\text{prep}}$ and $T$-basis measurements of $\tilde{S}^{\mathcal{G}}_{\text{consump}}$ on the blue memory qubits via the yellow and light yellow ancillary bus patches. Therefore, the buses must be connected, continuous, and have clear access to memory patches as the schedules require. The bilinear memory register and ancilla bus pattern must be at least three tiles wide to enable its comb-like pattern, allowing complete graph state processing. The logical width of the bilinear pattern per module (in the direction perpendicular to the boundaries of the bilinear pattern and $T$-transfer bus) is given by
$\begin{split}l_{\text{qbus}}=\max\left(\left\lceil\frac{\left\lceil\frac{{\tilde{n}}_{\text{logical}}}{n_{\text{per-leg}}^{\text{modules}}}\right\rceil}{2\left\lfloor\frac{l_{\text{edge}}-2}{4}\right\rfloor+1}\right\rceil,3\right)~{}.\end{split}$ (2)
The unallocated space, blank patches in fig. 6, are not accounted for by any explicit assignment but can still serve as ancillary buses for $O(1)$ operations.
Depending on the required accuracies for the complete algorithm and other architectural configurations, we select $n_{\text{distilleries}}$ copies of a suitable MSD factory ($T$-distilleries) per module to generate $T$-states at all steps continuously for consumption operations as needed - see also [11, 19] and section I.4.6. We lay out the factories on the opposite side of the bilinear pattern. The number of factories per module is
$\begin{split}n_{\text{distilleries}}=\lfloor\frac{l_{\text{edge}}-1}{\lceil\frac{L_{\text{width}}}{\sqrt{2}d}\rceil}\rfloor n^{\text{col}}_{\text{distill}}\geq 1~{},\end{split}$ (3)
and each have a physical size $L_{\text{length}}\times L_{\text{width}}$. Furthermore, the number of distillation columns (purple in fig. 6) in each module is given by
$\begin{split}n^{\text{col}}_{\text{distill}}=\lfloor\frac{l_{\text{edge}}-l_{\text{qbus}}-1}{\lceil\frac{L_{\text{length}}}{\sqrt{2}d}\rceil+1}\rfloor~{}.\end{split}$ (4)
The quantum memory and ancilla portions also serve as the endpoint for graph consumption. We do not allow the teleportation of $T$-states between modules. During a graph consumption step, the local generation, queue, and teleportation of $T$-states continue to facilitate all $T$ and $R_{z}$-basis measurements on graph-state nodes corresponding to the subschedule operations. We designed a specific mid-way ancillary system called a $T$-transfer bus or buffer (teal in fig. 6) to streamline this process. Once we allocate the quantum memory and ancilla portions, we can arrange the $T$ transfer bus in another comb-like pattern wrapped around $T$-factories to provide a continuous interface and maximize $n_{\text{distilleries}}$. The T-transfer bus is responsible for actively queuing the magic states, distilled by $n_{\text{distilleries}}$ factories, up to its logical length given by
$\begin{split}l_{\text{transfer-bus}}=&\left(l_{\text{edge}}-l_{\text{qbus}}-n^{\text{col}}_{\text{distill}}\left\lceil\frac{L_{\text{length}}}{\sqrt{2}d}\right\rceil\right)l_{\text{edge}}\\ &+n^{\text{col}}_{\text{distill}}\left\lceil\frac{L_{\text{length}}}{\sqrt{2}d}\right\rceil~{}.\end{split}$ (5)
The $T$-transfer bus incorporates a central $T$-buffer area as depicted in fig. 6. This buffer connects to the ancillae of the bilinear pattern along one entire edge of the module. In the simplest form of $T$-routing for graph consumption, the architecture ensures that the $T$-transfer bus can concurrently service at most $n^{\prime}=\min(n_{\text{distilleries}},n^{\text{row}}_{\text{qbus}})$ nodes of the ancilla bus with $T$-states through $O(1)$ operations (this should be considered a conservative lower limit for the number of simultaneous magic states accessible). Here, $n^{\text{row}}_{\text{qbus}}=\lfloor\frac{\left\lceil{\tilde{n}}_{\text{logical}}/n_{\text{per-leg}}^{\text{modules}}\right\rceil+1}{l_{\text{qbus}}}\rfloor$. Note that there can be at most $\left\lceil{\tilde{n}}_{\text{logical}}/n_{\text{per-leg}}^{\text{modules}}\right\rceil$ nodes requiring $T$-basis measurements simultaneously at any given time. Upon allocating all essential components, there may be none or many partially or fully unallocated columns of logical qubits left on the right side; fig. 6 shows a single leftover column as an example.
I.3.3 Computation of hardware’s thermal resources
Besides space-time volume data, including algorithmic execution times, physical space, and logical qubit counts, RRE can also identify the hardware resources required to execute specific algorithms. These include the counts of signal lines, the functional chip area, the number of amplifiers, etc. Heat loads pose a significant concern for superconducting qubits that operate at milliKelvin temperatures. Thus, we also estimate the power consumption and energy needed to run the FTQC.
Cryogenic energy consumption calculations for the proposed architecture are based on per-line thermal loads. These loads are calculated assuming a signaling solution with all-superconducting input and output lines below 4 K, conventional dispersive readout using a Traveling Wave Parametric Amplifier (TWPA), and High-Electron-Mobility Transistors (HEMT). We assume each qubit requires a microwave line for XY control and a flux bias line for tuning the qubit frequency. Each pair of qubits includes a tunable coupler [65] that requires a single flux line. Dispersive readout is expected to read out 10 qubits multiplexed in parallel. Heat loads are currently dominated by HEMT dissipation and static heat loads, so only those are considered. Example numbers are provided in RRE based on estimates from [66]. The power consumed by the overall system is scaled by rough estimates for the power necessary to operate 4 K cryo-plants [67] (500 kW of power for 500 W of cooling at 4 K) and dilution refrigerators (1 kW of power for 1 $\mu$W of cooling at 20 mK), as outlined in published work [67] and manufacturer specifications. We anticipate these estimates to evolve as they depend on the detailed designs of the signal chain.
I.4 Modular graph-state-based compilation and estimations
[…]
$\begin{split}\Big{|}\text{output}\Big{\rangle}=&U\ket{\text{input}:=s_{0}=0,s_{1}=0,\cdots,s_{n_{\text{input}}-1}=0}.\end{split}$ (6)
In practice, $U$ consists of a sequence of quantum instructions that may include tens of millions of logical gates. We always start with an all-zero |$\ket{\text{input}}$ state to simplify calculations.
One can always transpile the initial circuit to express it solely in terms of logical Cliffords, $T$, $T^{\dagger}$, and arbitrary-angle non-Clifford $R_{Z}$ gates, i.e. $U\big{(}\{\text{Clifford},T,T^{\dagger},R_{z}(\theta)\}\big{)}$. We denote the exact number of logical $T$ and $T^{\dagger}$ gates, excluding those implicitly included within the arbitrary-angle rotations, as $n^{T}_{\text{init}}$. Moreover, $n^{R_{z}}_{\text{init}}$ represents the exact number of arbitrary-angle non-Clifford $R_{Z}$ gates in $U$, considering those implicitly included in the logical gates (exclude $T$ and $T^{\dagger}$ operations counted separately in $n^{T}_{\text{init}}$). We assume fault-tolerant execution of $U\big{(}\{\text{Clifford},T,T^{\dagger},R_{z}(\theta)\}\big{)}$ using rotated surface code patches, as described in [12, 17, 18], with a code distance $d$. In our approach, the decomposition of non-Clifford rotations takes place after compilation during the consumption stages of the graphs (see fig. 2). This process is sometimes referred to as gate synthesis [68, 69]. We denote the diamond norm precision for gate syntheses as $\varepsilon$.
We consider an MBQC framework that employs ASG compilation for circuit execution. One benefit of this method is that it does not necessitate initializing the entire graph state, $\mathcal{G}$, on FTQC at all times [33, 70]. Instead, we divide circuit execution into several time steps, each containing a subcircuit or widget with its own (sub)graph. We can designate the parts of the graph state needed at any given moment by constructing three full schedules through sequentially appending sub-schedules across time steps: a preparation schedule for the graph state, labeled as $\tilde{S}^{\mathcal{G}}_{\text{prep}}$, which directs how to initialize $\mathcal{G}$, corresponding to the quantum state, based on the original entangling operations and other requirements for each step [34, 44]. A consumption schedule for the graph state, labeled as $\tilde{S}^{\mathcal{G}}_{\text{consump}}$, which specifies how to perform logical $R_{Z}(\theta)$-basis measurements relevant to executing the original non-Cliffords [33, 45, 71]. Additionally, a related schedule will identify the measurement bases used by FTQC during the consumption stage, which we refer to as $\tilde{S}^{\mathcal{G}}_{\text{meas}}$.
We verify that the widgetized schedules $\{\tilde{S}^{\mathcal{G}}_{\text{prep}},\tilde{S}^{\mathcal{G}}_{\text{consump}},\tilde{S}^{\mathcal{G}}_{\text{meas}}\}$ offer a comprehensive representation of the preparation and consumption of $\mathcal{G}$, the quantum state, in accordance to $U$. This is similar to any valid set of $\{S^{\mathcal{G}}_{\text{prep}},S^{\mathcal{G}}_{\text{consump}},S^{\mathcal{G}}_{\text{meas}}\}$ when time-sliced widgetization is not utilized. Previous studies [31, 12, 34, 44, 33, 45, 71] detail how valid arrangements of such schedules can effectively prepare [34] and consume [33] the graph, highlighting their connection to its spatial and temporal structure. Our focus will be on introducing the notation for $\{\tilde{S}^{\mathcal{G}}_{\text{prep}},\tilde{S}^{\mathcal{G}}_{\text{consump}},\tilde{S}^{\mathcal{G}}_{\text{meas}}\}$.
$\tilde{S}^{\mathcal{G}}_{\text{prep}}$ is a time-ordered list of $n_{\text{widgets}}-1$ sets, each acting as the preparation sub-schedule for a corresponding widget. The FTQC must follow these sub-schedules to initialize subgraphs in several sub-steps using $n$-body $CZ$ and single-qubit Clifford gates [31, 34, 44] corresponding to lattice surgery’s multi-product Pauli measurements of $X$ and $Z$-stabilizers [25, 26, 27, 28, 12]. Specifically, $\tilde{S}^{\mathcal{G}}_{\text{prep}}=\{\tilde{S}^{g_{0}}_{\text{prep}},\dots,\tilde{S}^{g_{n_{\text{widgets}}-1}}_{\text{prep}}\}$. Each preparation sub-schedule may comprise multiple sub-steps, represented as $\tilde{S}^{g_{i}}_{\text{prep}}=\{\tilde{S}^{g^{0}_{i}}_{\text{prep}},\dots,\tilde{S}^{g^{J(i)}_{i}}_{\text{prep}}\}~{}\forall~{}i\in\{0,\dots,n_{\text{widgets}}-1\}$. Furthermore, a $\tilde{S}^{g^{j}_{i}}_{\text{prep}}$ can contain several non-overlapping multi-product Pauli measurements [34, 44]. The multi-product Pauli measurements within a sub-step can be executed simultaneously using the ancillary buses.
Similarly, $\tilde{S}^{\mathcal{G}}_{\text{consump}}$ is a sequence of $n_{\text{widgets}}-1$ sets, with each set functioning as the consumption sub-schedule of a corresponding widget. To implement the original non-Cliffords, the FTQC must follow these sub-schedules, performing logical arbitrary-angle $R_{Z}(\theta)$-basis measurements on the relevant nodes in multiple sub-steps. In detail, $\tilde{S}^{\mathcal{G}}_{\text{consump}}=\{\tilde{S}^{g_{0}}_{\text{consump}},\dots,\tilde{S}^{g_{n_{\text{widgets}}-1}}_{\text{consump}}\}$. Each of these consumption sub-schedules may contain several sub-steps, represented as $\tilde{S}^{g_{i}}_{\text{consump}}=\{\tilde{S}^{g^{0}_{i}}_{\text{consump}},\dots,\tilde{S}^{g^{K(i)}_{i}}_{\text{consump}}\}~{}\forall~{}i\in\{0,\dots,n_{\text{widgets}}-1\}$. A specific $\tilde{S}^{g^{k}_{i}}_{\text{consump}}$ can contain multiple non-overlapping measurement lists detailed in [33, 71]. The measurements found in the non-overlapping lists within a sub-step can be executed simultaneously. Meanwhile, $\tilde{S}_{\text{meas}}^{G}$ serves as a connected sequence that details the rotational basis ($T$ or $R_{Z}(\theta)$) required for node measurements across the corresponding sub-steps of $\tilde{S}^{\mathcal{G}}_{\text{consump}}$. This method differentiates between two categories of node measurements during the consumption stage: those that need $T$-basis measurements consuming a single distilled $T$ and those that require $R_{Z}(\theta)$-basis measurements consuming multiple distilled $T$ states following gate-synthesis of $R_{Z}(\theta)$.
The number of nodes in $\mathcal{G}$, graph size $N$, is typically much larger than the number of input qubits, $N\gg n_{\text{input}}$. This is because we always add ancillary qubits for an MBQC implementation of a given quantum algorithm ($N$ is always greater than $n_{\text{input}}$ [71, 32, 70]). However, due to the scheduling requirements above, only a relatively small subset of logical qubits, each representing a graph node, is required at any given step to execute the unitary operation $U$. We denote the maximum number of logical qubits needed at any step as $\tilde{n}^{\mathcal{G}}_{\text{logical}}$, i.e., the maximum quantum memory required. Typically, $n_{\text{input}}<\tilde{n}^{\mathcal{G}}_{\text{logical}}<N$ [71, 33].
One can create distinct graph states with similar attributes from different input states, but the same $U$. We construct $\mathcal{G}$ by stitching graphs from each time step for resource estimation purposes. To admit arbitrary input nodes into these ASGs, we always initialize all nodes in the all-zero input state, |$\ket{0,0,\dots,0}$, and then replace the input nodes with $n_{\text{input}}$ copy of (|$(\ket{0+}+\ket{1-})^{\otimes n_{\text{input}}}$ attaching the appropriate ancillary nodes. This process allows us to perform teleportation from the output nodes of one (sub)graph to the input nodes of another belonging to the consecutive time step [71, 32], which we utilize to verify and stitch graph states.
To summarize, we can achieve exact resource estimations of $U$ by providing an MBQC-based set of primitives for the non-widgetized algorithm:
$\begin{split}S_{\text{est}}&(U,\{\text{config}\}):=\\ &\{d,\varepsilon,n^{T}_{\text{init}},n^{R_{z}}_{\text{init}},n^{\mathcal{G}}_{\text{logical}},S^{\mathcal{G}}_{\text{prep}},S^{\mathcal{G}}_{\text{consump}},S^{\mathcal{G}}_{\text{meas}}\},\end{split}$ (7)
In this equation, the required user inputs are explicitly stated on the left side, while the right side generates the corresponding set of primitives needed for resource estimation. It becomes incredibly challenging to construct $S_{\text{est}}$ for a utility-scale $U$ containing tens of millions of gates and considerable circuit depth, whether for resource estimation or hardware execution purposes. Therefore, we proposed a time-sliced widgetization of $U$ leading to an approximate resource estimation set $\tilde{S}_{\text{est}}$. Below, we provide a step-by-step recipe to create and rigorously verify $\tilde{S}_{\text{est}}$.
I.4.1 Input circuit widgetization
Figure 7: Input circuit decomposition strategy: This visualization demonstrates how we "widgetize" a complete logical algorithm, $U$, into $n_{\text{widgets}}$ time-sliced widgets represented as $U_{n_{\text{widgets}}-1}\dots U_{1}U_{0}$. Each widget has the same width, $n_{\text{input}}$, but varied depth. Some widgets may be repeated in the time direction (for example, $U_{3}\equiv U_{0}$ above), allowing us to compile only $n^{*}_{\text{widgets}}$ distinct subcircuits.
Due to the size of a utility-scale input circuit, a decomposition strategy is necessary to prevent memory bottlenecks on RRE. Fortunately, quantum algorithms typically consist of repeating blocks, aiding their decomposition.
Our decomposition strategy involves slicing the unitary operator $U$ in the time direction into $n_{\text{widgets}}$ widgets, or subcircuits, of fixed width $n_{\text{input}}$. Each widget may have a distinct and emergent depth. It is essential to emphasize that no entangling gates connect adjacent subcircuits. We can then execute the complete algorithm as a sequence of these subcircuits, expressed as $U=U_{n_{\text{widgets}}-1}\dots U_{1}U_{0}$, where
$\begin{split}U_{i}\ket{\text{state}}_{i}=\ket{\text{state}}_{i+1}~{}\text{for}~{}i\in\{0,\cdots,n_{\text{widgets}}-1\}~{}.\end{split}$ (8)
Here, |$\ket{\text{state}}_{0}\equiv\ket{\text{input}}$ and |$\ket{\text{state}}_{n_{\text{widgets}}}\equiv\ket{\text{output}}$.
In a typical quantum algorithm, many widgets may be equivalent and repeated throughout the time direction. In such cases, we only need to compile the set of distinct widgets, which can be, for example, $[U_{0},U_{4},\dots,U_{n_{\text{widgets}}-7}]$. We use $n^{*}_{\text{widgets}}$ to represent the number of distinct widgets in the set $[U_{0},U_{1},\dots,U_{n_{\text{widgets}}-1}]$ ($n^{*}_{\text{widgets}}\leq n_{\text{widgets}}$). We refer to this process as circuit widgetization, illustrated in fig. 7. We based our particular choice of widgetization strategy on creating subcircuit dependency graphs. This approach allows us to compile enormous circuits into elementary operations and unitarily reconnect compiled widgets, avoiding excessive memory overheads. Compilation is essential for executing the algorithm and accurate resource estimations. Additionally, this strategy helps to keep the number of distinct widgets $n^{*}_{\text{widgets}}$ relatively low, leads to ease of implementation, and justifies using time slicing with a fixed $n_{\text{input}}$. However, this particular choice of widgetization is likely not optimized to minimize space-time costs. We present further details on our widgetization strategy in section A.2.4.
I.4.2 Stitching strategy
Following the widgetization process, we can construct an approximate version of $S_{\text{est}}$. We propose the widgetization-based set $\tilde{S}_{\text{est}}$ as a good approximation of $S_{\text{est}}$ for resource estimation purposes, defined as follows:
$\begin{split}\tilde{S}_{\text{est}}&:=\\ &\{d,\varepsilon,n^{T}_{\text{init}},n^{R_{z}}_{\text{init}},\tilde{{n}}^{\mathcal{G}}_{\text{logical}},\tilde{S}^{\mathcal{G}}_{\text{prep}},\tilde{S}^{\mathcal{G}}_{\text{consump}},\tilde{S}^{\mathcal{G}}_{\text{meas}}\}\\ &\approx S_{\text{est}}~{}.\end{split}$ (9)
Recall that $n^{T}_{\text{init}}$ and $n^{R_{Z}}_{\text{init}}$ are the exact number of explicit $\{T,T^{\dagger}\}$ and arbitrary-angle $R_{Z}$ operations in $U$. In this widgetization-based estimation procedure, we first obtain estimation primitives for each widget based on its graph, sometimes referred to as a subgraph. We can find the overall circuit resource estimates as follows:
$\begin{split}&\tilde{n}_{\text{logical}}^{\mathcal{G}}:=\max_{g_{i}}(n^{g_{i}}_{\text{logical}})\approx n_{\text{logical}}^{\mathcal{G}},\\ &\tilde{S}^{\mathcal{G}}_{\text{prep}}:=\cup_{i=0}^{n_{\text{widgets}}-1}S^{g_{i}}_{\text{prep}}\equiv S^{\mathcal{G}}_{\text{prep}},\\ &\tilde{S}^{\mathcal{G}}_{\text{consump}}:=\cup_{i=0}^{n_{\text{widgets}}-1}S^{g_{i}}_{\text{consump}}\equiv S^{\mathcal{G}}_{\text{consump}},\\ &\tilde{S}^{\mathcal{G}}_{\text{meas}}:=\cup_{i=0}^{n_{\text{widgets}}-1}S^{g_{i}}_{\text{meas}}\equiv S^{\mathcal{G}}_{\text{meas}}~{}.\end{split}$ (10)
Here, $g_{i}$ represents the graph state compiled from the $i$th widget.
Equation 10 emphasizes the key assumptions we made to generate our estimates, formally outlining our stitching strategy. I.e.  how we combine the individual graphs and scheduling lists to approximate the complete $\mathcal{G}$, as well as $S^{\mathcal{G}}_{\text{prep}}$, $S^{\mathcal{G}}_{\text{consump}}$, and $S^{\mathcal{G}}_{\text{meas}}$. While an FTQC needs to physically stitch the widgets (through teleportations) to execute the complete circuit, the same is unnecessary for resource estimation purposes. Consequently, we reduced the computational demands for obtaining resource estimates by calculating the widgetized quantities in eq. 10 through software.
I.4.3 Pre-processing the input subcircuits
To prepare for graph compilation, we must parse each widget and transpile it to include only logical Clifford+$T$+$R_{Z}(\theta)$ gates, using exact gate equivalences. Here, the $\theta$ values represent non-Clifford angles. Later, during the consumption stage, when we want to perform single-qubit logical measurements, we need to further decompose each $R_{Z}(\theta)$ node into a sequence of Clifford+$T$ gates through gate synthesis [68, 69].
[…]
Once transpilation is complete, the next step is to generate the unified MBQC-based set of graph states $\{g\}$ and schedules $\{S^{g}_{\text{prep}}\}$, $\{S^{g}_{\text{consump}}\}$, and $\{S^{g}_{\text{meas}}\}$ (and other essential intermediary lists detailed below), which leads to $\tilde{S}_{\text{est}}$ sufficient for the resource estimations of $U$.
I.4.4 Cabaliser: compiling to graphs and consumption schedules
RRE utilizes the stabilizer tableau-based fault-tolerant compiler Cabaliser [71] as a crucial intermediate layer to transform the set of individual logical widgets, $\{U\}$, into corresponding MBQC-based subgraphs, consumption schedules, “locally simulated” quantum states, and other essential information. Locally simulated quantum states are Cabaliser’s outputs derived through exact calculations. Cabaliser effectively maps an input state |$\ket{\text{state}}_{i-2}$ into |$\ket{\text{state}}^{\text{Cabaliser}}_{i}$: it applies $U_{i}$, Pauli corrections, the consumption subschedule, and additional logical operations (such as Hadamards) to ($i-1$)th graph state to create a faithful vector representation of the corresponding universal subgraph at step $i$. It is guaranteed that |$\ket{\text{output}}\equiv\ket{\text{state}}^{\text{Cabaliser}}_{n_{\text{widget}}-1}$. This vector representation helps verify the unitaritiness of RRE and Cabaliser’s outputs for small $n_{\text{input}}$. We present a systematic method to verify the graphs and schedules produced by RRE, taking into account widgetization and stitching in section A.3. For complete details on scalability and various compilation components of Cabaliser, we refer readers to [71, 38, 31].
For a given individual subcircuit $U_{i}$, which includes only Clifford+$T$+$R_{Z}(\theta)$, Cabaliser can generate the following set of quantities:
[…]
In this equation, $S_{\text{frames}}^{g_{i}}$ is a set that tracks the required Pauli corrections (frames) to be applied to the measurements at the end. $S_{\text{output}}^{g_{i}}$ and $S_{\text{input}}^{g_{i}}$ represent the sets (or maps) of output and input nodes in the graph state, respectively. Since we decompose the entire algorithm in the time direction with a uniform width that covers all qubits, the number of intermediate input and output nodes remains the same. $n^{g_{i}}_{\text{logical}}$ denotes the quantum memory needed to process $U_{i}$. $S_{\text{consump}}^{g_{i}}$ identifies the graph consumption schedule for performing individual logical qubit measurements. $S_{\text{meas}}^{g_{i}}$ identifies which rotational basis we must use for each measurement during the consumption stage. Finally, |$\ket{\text{state}}^{\text{Cabaliser}}_{i}$ represents the locally simulated quantum state extracted from the outputs generated up to $i$th step.
We repeat the process of eq. 11 for all $n^{*}_{\text{widgets}}$ distinct widgets, providing the user with the information required for the approximate estimation set in eq. 9, excluding the preparation schedules, $\{S^{g}_{\text{prep}}\}$, which we generate using another tool detailed below.
I.4.5 Substrate Scheduler: generating preparation schedules
RRE utilizes the Substrate Scheduler [44] to map the graph states $\{g\}$, calculated by Cabiliser in eq. 11, to a set of graph preparation sub-schedules $\{S_{\text{prep}}^{g}\}$. These preparation sub-schedules are optimized for the quantum memory register and ancilla bus in the superconducting hardware layout described in fig. 1.
For the Substrate Scheduler calculations, we assume a dual-rail or bilinear quantum memory and ancilla bus pattern consisting of linear arrangements of surface code qubits for storing graphs and mediating logical measurements on the nodes as per fig. 1. The pattern has a fixed size $\tilde{n}_{\text{logical}}:=\max_{g_{i}}(n^{g_{i}}_{\text{logical}})$ per rail. This assumption ensures no memory constraints when executing the preparation and consumption stages consecutively. While the intra-module architecture we use for resource estimates follows fig. 6, the Substrate Scheduler results can still be matched and utilized for the wrapped bilinear pattern. Currently, the Substrate Scheduler identifies an optimal schedule based on the individual mapping of each $g$ to this layout. The Substrate Scheduler’s performance could be enhanced by incorporating additional information about node types and supporting customized bus architectures, which we plan to address in future work.
I.4.6 Space-time-optimal resource estimation
DistillWidget (superconducting) $p_{out}$ Phys. Space($L_{\text{width}}\times L_{\text{length}}$) Qubits, $Q~{}$ Cycles, $C$
(15-to-1)17,7,7 $4.5\times 10^{-8}$ 64$\times$72 4620 42.6
[…]
Table 1: Distillation Widget Lookup Table: Litinski’s distillation widgets that match our superconducting architecture with $p=10^{-3}$. Extended from [11, Table 1].
[…]
At a macroscopic level, the architecture features a ladder-style structure with $n_{\text{per-leg}}^{\text{modules}}$ pairs of QPU modules connecting the legs, as illustrated in fig. 3. We based this design on the assumption that superconducting platforms have a low syndrome extraction cycle and would benefit from minimizing the physical footprint while interleaving graph preparation and consumption. Consequently, we always create $\tilde{n}_{\text{logical}}$ surface code patches for logical memory patches shared among all modules on a leg of the macro-architecture.
At the intra-module level, we organize each module as a square grid of logical qubits (patches), the foundational building blocks for all components. The edge size is determined by $l_{\text{edge}}=\lfloor\sqrt{\frac{N_{\text{phys}}}{2d^{2}}}\rfloor$, where $N_{\text{phys}}=1,000,000$ represents the total number of physical qubits available per module. We assumed each rotated surface code patch contains $2d^{2}$ physical qubits.
We performed a static allocation of intra-modular components, providing sufficient resources for the largest widgets as illustrated in fig. 6. Each module has five main components to which we allocate physical qubits. The unallocated or "wasted" space can still be utilized for some ancillary operations. The first component is a quantum register comprising memory or graph-states qubits. The second is a chain of ancilla bus qubits of the same length, facilitating logical measurements on the memory qubits, which connects to the third component, i.e., the supplementary bus. The first two components are arranged in a snake-like pattern to occupy one QPU edge fully. The fourth component is the linear $T$-transfer bus to supply the ancilla bus portion with $T$-states for graph consumption. The last is $T$-distilleries, also called distillation widgets or factories. See section I.3.2 for details.
The space-time volume of the complete graph creation on the bi-linear quantum bus before single-qubit logical measurements is
$\begin{split}V_{\mathcal{G}}=2\tilde{n}_{\text{logical}}\mathbb{L}_{\text{prep}}d,\end{split}$ (12)
where $\mathbb{L}_{\text{prep}}(\tilde{S}_{\text{prep}})=\sum_{\text{sub}\in\tilde{S}_{\text{prep}}}L_{\text{sub}}$. Here, $L_{\text{sub}}$ is the number of distinct measurement steps in the preparation sub-schedule of sub. In other words, $\mathbb{L}_{\text{prep}}$ is the total number of distinct measurement steps required for the complete graph preparation.
[…]
In “Distillation Widget Lookup Table”, table 1, the output probability, denoted as $p_{\text{out}}$, determines the error rate for the ancillary state of factories and sets an upper limit on the error rate for any $T$-gate implemented using the ancilla. The specifications for the number of qubits, the dimensions of the 2D patch, and the number of physical cycle sweeps all depend on the construction of the factory. Our analysis focused exclusively on selecting distillation widgets that produce $T$-states, thus excluding any entries related to $CCZ$ from the original [11, Table 1].
Next, we must include the resource cost for decomposing non-Clifford into Clifford+$T$ (gate synthesis) to consume graph-state nodes through single-qubit measurements. After completing a graph preparation sub-schedule, at each sub-step of the consumption schedule, we utilize one or more purified $T$-resources to perform measurements on the desired basis. The FTQC must teleport the $T$-states from distillation factories to the ancilla bus. We assume that the graph nodes are either of the type that requires arbitrary $R_{Z}$-basis measurements or a single $T$-basis measurement. Measurements that involve $R_{Z}$-axis rotations are decomposed into sequences of $T$-gates, with a fixed worst-case length of $L_{\varepsilon}=\lceil c_{0}\log(1/\varepsilon)+c_{1}\rceil$. We determine this length through the gate synthesis methods available to RRE. Here, $\varepsilon$ refers to the diamond-norm error for decompositions into Clifford+$T$—see [68, 69] for further details.
The generation of each distilled $T$-state in the factories requires $C$ cycles, and we need one further surface code tock (equivalent to $d$ cycles) to interact it with the graph node and measure it in either the $X$ or $Z$ basis. In this work, we ignore the time cost of any operation that requires $O(1)$ cycles (e.g., shuffling surface code patches), assuming we can perform them simultaneously during sub-steps at no additional cost. Therefore, for each non-Clifford measurement of memory nodes, assuming an $R_{Z}$-basis as an example, the volume increases as follows:
$\begin{split}2\tilde{n}_{\text{logical}}\mathbb{L}_{\text{prep}}d\rightarrow&2\tilde{n}_{\text{logical}}\mathbb{L}_{\text{prep}}d+\\ &(2\tilde{n}_{\text{logical}}+n_{\text{per-leg}}^{\text{modules}}l_{\text{transfer-bus}})(C+d)L_{\varepsilon}~{}.\end{split}$ (13)
In eq. 13, the second term reflects the fact that both graph node consumption and distillation occur over a space of $2\tilde{n}_{\text{logical}}+n_{\text{per-leg}}^{\text{modules}}l_{\text{transfer-bus}}$. All nodes with distinct measurement basis types, $T$ and $R_{Z}$, must be measured sequentially.
If we assume the worst-case sequential execution, the upper bound for full space-time volume becomes:
$\begin{split}V^{\text{sequential}}_{\text{total}}&=2\tilde{n}_{\text{logical}}\mathbb{L}_{\text{prep}}d+\\ &(2\tilde{n}_{\text{logical}}+l_{\text{transfer-bus}})(N^{\text{seq}}_{\text{consump}}d+N^{\text{seq}}_{\text{distill}}C)~{}.\end{split}$ (14)
We define the total number of sequential T-states for graph consumption as $N^{\text{seq}}_{\text{consump}}=\left\lceil\frac{n^{T}_{\text{init}}}{n^{\prime}}\right\rceil+L_{\varepsilon}\left\lceil\frac{n^{R_{z}}_{\text{init}}}{n^{\prime}}\right\rceil$. Recall that we assumed a simplified $T$-routing scheme that allows the ancilla bus to serve up to $n^{\prime}$ simultaneous magic states. FTQC can only consume distinct $T$ and $R_{Z}$ nodes in parallel. The number of sequential T-states to distill is $N^{\text{seq}}_{\text{distill}}=\left\lceil\frac{n^{T}_{tot}}{n^{\prime}}\right\rceil$, where the total $T$-state measurements required is defined as $n^{T}_{tot}:=n^{R_{Z}}_{\text{init}}L_{\varepsilon}+n^{T}_{\text{init}}$. The quantity $n^{T}_{tot}$ is the proxy we use for $T$-count and typically becomes larger than $N_{\mathcal{G}}$. We do not include distillation widget space-time volumes, as their errors subsume into the output probability $p_{out}$ of the distilled state [11, Fig. 1]. If the select widget from table 1 has a second (20-to-4) layer (the second and third widgets), then we must replace $n^{\prime}\rightarrow 4n^{\prime}$ in all space-time equations.
As detailed in section I.3.1, a single cycle time of the surface code, $p_{C}$, has a failure probability characterized by the physical and logical error rates relationship given in eq. 1. Consequently, we can express the failure probability of a single unit cell containing $d$ cycles as:
$p_{\text{log-cell}}=1-(1-p_{C})^{d}.$ (15)
The failure rate for a total volume of these cells, denoted as $V^{\text{sequential}}_{\text{total}}$, must satisfy the following condition:
$p_{\text{algo-fail}}>1-(1-p_{\text{log-cell}})^{V^{\text{sequential}}_{\text{total}}},$
[…]
Assuming that $p_{C}\ll 1$, we can approximate $(1-p_{C})^{d}\approx 1-p_{C}d$, leading us to conclude that $p_{\text{log-cell}}\approx p_{C}d$ and consequently,
$p_{\text{algo-fail}}\gtrsim 1-\left(1-p_{C}d\right)^{V^{\text{sequential}}_{\text{total}}}.$
This equation highlights the relationship between the failure rates and the characteristic space-time volume for the QEC process.
Subtracting each side of section I.4.6 from 1 and taking the logarithm of both sides, we get
$\log\left(1-p_{\text{algo-fail}}\right)\lesssim$ $V^{\text{sequential}}_{\text{total}}\log\left(1-p_{C}d\right)$
$\lesssim$ $-V^{\text{sequential}}_{\text{total}}p_{C}d.$ (16)
If we substitute $V^{\text{sequential}}_{\text{total}}$ from eq. 14 into eq. 16, the primary logarithmic inequality to solve can be expressed as
$\begin{split}\kappa d&\left(\frac{p}{p_{\text{thresh}}}\right)^{(d~{}+1)/2}\big{(}2\tilde{n}_{\text{logical}}\mathbb{L}_{\text{prep}}d~{}+(2\tilde{n}_{\text{logical}}\\ &+n_{\text{per-leg}}^{\text{modules}}l_{\text{transfer-bus}})(N^{\text{seq}}_{\text{consump}}d+N^{\text{seq}}_{\text{distill}}C)\big{)}<-J_{1}\end{split}$ (17)
where $J_{1}=\log(1-p_{\text{algo-fail}})$. The minimum code distance required is the smallest integer $d$ that satisfies eq. 17. The only free variables are the cycle time of the distillation widget, $C$, and $\varepsilon$. We previously extracted $\mathbb{L}_{\text{prep}}$ and $\tilde{n}_{\text{logical}}$ from the Cabaliser and Substrate Scheduler.
To determine $C$ and $\varepsilon$, we perform nested loops to iterate through widgets and then gate synthesis’ $\varepsilon$. The inner loop feeds back into the Clifford+$T$ decomposition as the failure rate of a unit cell, $p_{\text{log-cell}}$, establishes the approximation error as $\varepsilon<p_{\text{log-cell}}$. By selecting a specific widget, we fix $C$, which also determines the output error rate of the distilled state, $p_{\text{out}}$.
In a compiled quantum algorithm, we aim to ensure that the distilled $T$-gates have precisely the same error rates as any other gates in the compiled circuit. A Pauli initialization or measurement occupies a single unit cell in the space-time volume. Thus, its failure rate is governed by eq. 15. This error rate must be dominant. Therefore, the error output, $p_{out}$, from the distillation circuit should be less than the failure rate of this cell (that is, we must ensure that the distillation output is at least as good as a unit cell in the space-time volume). Consequently, we first select the smallest value of $C$ from table 1 and solve eq. 17 for $d$. We then verify that, for a physical error rate of $p=10^{-3}$ and our calculated solution for $d$, whether
$p_{\text{log-cell}}=1-(1-p_{C})^{d}>p_{\text{out}}$ (18)
does hold for the corresponding $p_{\text{out}}$ entry. If this condition is unmet, we must choose a better-performing $T$-distillation widget in the outer loop. We then proceed to the next entry on the list and repeat until we identify a distillation widget that performs at least as well as the failure of a logical unit cell.
Once a specific distillation widget is selected, the total number of allocated physical qubits across all modules at any given time and the total intra-modular runtime (including the first subgraph preparation time at step-00, all consumption operations, and delays) can be calculated as follows:
$\begin{split}N&{}_{\text{alloc-phys}}=\\ &4\tilde{n}_{\text{logical}}d^{2}+n^{\text{modules}}_{\text{per-leg}}(n_{\text{distilleries}}Q+2d^{2}l_{\text{transfer-bus}}),\\ t&{}^{\text{total}}_{\text{consump}}=\\ &8td(L^{0}_{\text{prep}}+N^{\text{seq}}_{\text{consump}})+t^{\text{delay}}_{\text{distill}}+t^{\text{delay}}_{\text{prep}}~{}.\end{split}$ (19)
In eq. 19, $n^{\text{modules}}_{\text{per-leg}}$ refers to the number of modules per leg of the ladder macro-architecture – that is, there are $2n^{\text{modules}}_{\text{per-leg}}$ modules in total. $t$ denotes the characteristic gate time, establishing the architecture’s quantum tock for all intra-modular operations. The factor of 8 arises from the fact that in our physical architecture, each surface code sweep for syndrome extraction circuits necessitates one initialization, two Hadamards, four entangling gates, and one measurement, all assumed to take a duration of $t$ (see, for instance, [10, Fig. 9]). The temporal cost of initializing the first subgraph $g_{0}$ with $L^{0}_{\text{prep}}(S^{0}_{\text{prep}})$ is always nonzero and was explicitly added. FTQC carries out the remaining graph preparations and $T$-state distillation operations in parallel with consumption operations. There may be overall delays accumulated over different time steps, defined as $t^{\text{delay}}_{\text{distill}}=\sum_{i}t^{\text{distill-delay}}_{\text{i}}$ and $t^{\text{delay}}_{\text{prep}}=\sum_{i}t^{\text{prep-delay}}_{\text{i}}$, which we detail below.
To execute each consumption sub-schedule $i$, a specific number of distilled magic states must exist in the module’s $T$-transfer bus. We must introduce a delay if we require more magic states than what was accumulated in the $T$-transfer bus during the prior preparation step, $i-1$. The number of $T$ states per fridge after a preparation step is
$\begin{split}n^{T-\text{per-fridge}}_{i}=\max\big{(}\lfloor&\frac{n_{\text{distilleries}}t^{\text{prep}}_{i}}{8tC}\rfloor,l_{\text{tansfer-bus}}\big{)}\\ &\text{for}~{}i\in\{0,\dots,n_{\text{widgets}}-1\}.\end{split}$ (20)
Here, we have defined the hybrid (intra- and inter-modular) time required to prepare the subgraph for preparation step $i$ as $t^{\text{prep}}_{i}=8d(N_{i}^{\text{intra-comms}}t+N_{i}^{\text{cross-fridges}}t_{\text{inter}})$. Where $N_{i}^{\text{intra-comms}}=\sum_{j\in L_{i,j}^{\text{prep}}}\lfloor\frac{\mathbb{L}_{i}^{\text{prep}}}{\sum_{k\in L_{i,j}^{\text{prep}}}\lfloor d^{\text{max}}_{i,j,k}/\tilde{n}_{\text{logical}}\rfloor}\rfloor~{}\text{for}~{}n^{\text{modules}}_{\text{per-leg}}>1$ is the count of intra-module operations, while $N_{i}^{\text{cross-fridge}}=\sum_{j\in L_{i}^{\text{prep}}}\lceil\frac{\sum_{k\in L_{i,j}^{\text{prep}}}\lfloor d^{\text{max}}_{i,j,k}/\tilde{n}_{\text{logical}}\rfloor}{n_{\text{inter-pipes}}}\rceil~{}\text{for}~{}n^{\text{modules}}_{\text{per-leg}}>1$ represents the number of cross-fridge operations in the preparation sub-schedule $i$ (only in the vertical direction along a ladder leg). We assume that $t_{\text{inter}}\gg t$ is the characteristic physical timescale for all inter-module communications. Moreover, we define $d^{\text{max}}_{i,j,k}$ as the maximum distance between any pairs of graph nodes in the tuple $k$ of sub-list $j$ of preparation sub-schedule $i$ (see [44] for details), and $n_{\text{inter-pipes}}$ is the number of inter-modular pipes connecting each pair of modules in the macro-architecture.
We can now define the distillation delay for step $i\in\{0,\dots,n_{\text{widgets}}-2\}$ as
$t^{\text{distill-delay}}_{\text{i}}=\begin{cases}8tC&\lceil\frac{L^{\text{max-seq}}_{\text{consump},i}-n^{T-\text{per-fridge}}_{i}}{n_{\text{distilleries}}}\rceil,\\ &\text{if }L^{\text{max-seq}}_{\text{consump},i}>n^{T-\text{per-fridge}}_{i}\\ 0,&\text{otherwise}\end{cases}$
where $L^{\text{max-seq}}_{\text{consump},i}=n^{\text{max}-T}_{i}+L_{\varepsilon}n^{\text{max}-R_{Z}}_{i}$ is the maximum sequential $T$-length per module, considering all consumption nodes of step $i$. We have defined $n^{\text{max}-T}_{i}=\max_{j}(n^{T}_{i,j})~{}\text{for}~{}i\in S^{\text{meas}}_{i},~{}j\in\{0,\dots,n^{\text{modules}}_{\text{per-leg}}-1\}$ as the maximum number of $T$-basis measurement nodes in all modules specified by measurement sub-schedule $S^{\text{meas}}_{i}$. Similarly, $n^{\text{max}-R_{Z}}_{i}=\max_{j}(n^{R_{z}}_{i,j})~{}\text{for}~{}i\in S^{\text{meas}}_{i},~{}j\in\{0,\dots,n^{\text{modules}}_{\text{per-leg}}-1\}$.
Similarly, for step $i$, an idle delay must be added if the preparation for the next subgraph $i+1$ is not yet complete (which runs parallel across the interleaving fridges in the other leg of the ladder). One can establish an upper bound for the individual time related to the intra-modular node consumption of step $i$ as $t^{\text{consump-intra}}_{i}=8td(\lceil\frac{n^{\text{max}-T}_{i}}{n_{\text{distilleries}}}\rceil+L_{\varepsilon}\lceil\frac{n^{\text{max}-Rz}_{i}}{n_{\text{distilleries}}}\rceil)$. Therefore, the preparation delay for step $i\in\{0,\dots,n_{\text{widgets}}-2\}$ becomes
$t^{\text{prep-delay}}_{\text{i}}=\begin{cases}t^{\text{prep}}_{i+1}&-t^{\text{consump-intra}}_{i}-t^{\text{distill-delay}}_{\text{i}},\\ &\text{if }t^{\text{prep}}_{i+1}>t^{\text{consump-intra}}_{i}+t^{\text{distill-delay}}_{\text{i}}\\ 0.&\text{otherwise}\end{cases}$
Lastly, we need to account for the total time required on the actual FTQC to stitch the output nodes of the graph at step $i$, located along one leg of the ladder, to the input nodes of the graph at step $i+1$, situated along the other leg of the ladder, for $i\neq n_{\text{widget}}-1$. We assume we can perform the hand-over process on-the-fly by teleporting data-qubit Bell pairs across the modules in a horizontal direction, which involves adding logical $CNOT$ or $CZ$ gates (or, for smaller problems, some equivalent $SWAP$ operations). We must add a cost if FTQC needs to teleport the graph nodes vertically across the same leg (infeasible through $O(1)$ operations). Therefore, an upper bound for this quantity can be expressed as
$\begin{split}t^{\text{total}}_{\text{handover-inter}}=8t_{\text{inter}}d\sum_{i\in\{0,\dots,n_{\text{widget}}-2\}}\lceil\frac{N^{\text{cross-module-IO}}_{i}}{n_{\text{inter-pipes}}}\rceil,\end{split}$ (21)
where we define the total number of cross-fridge communications needed to hand over output nodes of graph $i$ to $i+1$ as $N^{\text{cross-module-IO}}_{i}=\sum_{\tilde{o}\in S^{i}_{\text{output}},\tilde{i}\in S^{i+1}_{\text{input}}}n^{i}_{\tilde{o},\tilde{i}}$; here, $n^{i}_{\tilde{o},\tilde{i}}$ is the number of crossings necessary to hand over node $\tilde{o}$ of graph $i$ to node $\tilde{i}$ of graph $i+1$ for $i\neq n_{\text{widget}}-1$.
Finally, the total FTQC hardware time can be expressed as
$\begin{split}t_{\text{hardware}}^{total}=t^{\text{total}}_{\text{intra}}+t^{\text{total}}_{\text{handover-inter}}+t_{\text{decode-delay}}^{\text{total}}.\end{split}$ (22)
In this equation, $t_{\text{decode-delay}}^{\text{total}}$ represents the overall delay involved in performing decoding, entirely using classical methods, for all QEC cycles. We propose a method to estimate $t^{\text{decode-delay}}_{tot}$ in section A.1, based on generic decoders.
The following summarizes the workflow as Algorithm 1.
1
2
input : Logical quantum algorithm, QuantumCircuit, $U$
input : Look-up table: DistillationWidgets (Code, $C$, $Q$, $p_{\text{out}}$)
output : Total required physical qubits, $N_{\text{alloc-phys}}=4\tilde{n}_{\text{logical}}d^{2}+n^{\text{modules}}_{\text{per-leg}}(n_{\text{distilleries}}Q+2d^{2}l_{\text{tansfer-bus}})$
output : Total intra-modular time, $t^{\text{total}}_{\text{intra}}=8td(L^{0}_{\text{prep}}+N^{\text{seq}}_{\text{consump}})+t^{\text{delay}}_{\text{distill}}+t^{\text{delay}}_{\text{prep}}$
3
4$\tilde{n}_{\text{logical}}$, $\tilde{S}_{\text{consump}}$, $\tilde{S}_{\text{prep}}$, $\tilde{S}_{\text{meas}}$} = Cabaliser-Tracker-Scheduler(QuantumCircuit);
5
6for widget $\in$ DistillationWidgets do
7       if widget is (20-to-4) then
8            $n_{\text{distilleries}}\rightarrow 4n_{\text{distilleries}}$;
9      while $M_{\varepsilon}==\text{False}$ do
10             Solve $\kappa d\left(\frac{p}{p_{\text{thresh}}}\right)^{(d~{}+1)/2}\big{(}2\tilde{n}_{\text{logical}}\mathbb{L}_{\text{prep}}d~{}+(2\tilde{n}_{\text{logical}}+n_{\text{per-leg}}^{\text{modules}}l_{\text{transfer-bus}})(N^{\text{seq}}_{\text{consump}}d+N^{\text{seq}}_{\text{distill}}C)\big{)}<-J$ for the smallest integer, $d$, where $J=\log(1/(1-p_{\text{algo-fail}}))$, $p=10^{-3}$;
11             $p_{\text{log-cell}}=1-(1-\kappa(p/p_{\text{thresh}})^{(d+1)/2})^{d}$;
12             $M_{\varepsilon}=p_{\text{log-cell}}>\epsilon$;
13             if $M_{\varepsilon}==\text{False}$ then
14                  Decrease $\varepsilon$;
15
16      $M_{\text{log-cell}}=p_{\text{log-cell}}>p_{\text{out}}$;
17       if $M_{\text{log-cell}}==\text{True}$ then
18            Break loop;
19
20
Algorithm 1 RRE’s estimation methodology based on the middle-of-the-way architecture of fig. 2.
II Test cases
To demonstrate resource estimation using graph-state formalism, we generated circuits for quantum algorithms that target core computational challenges in real-world applications. Resource-efficient compilation necessarily means considering aspects of the hardware in the synthesis methodology and empirically evaluating alternatives when analytical results are not easily obtained. As such, these circuits are then compiled into fault-tolerant quantum computing patterns with specificity to superconducting architectures. This section summarizes the test cases chosen for circuit synthesis to demonstrate our resource estimation methodology.
II.1 Quantum Fourier Transform
The Quantum Fourier Transform (QFT) is a common component of quantum algorithms, including Quantum Phase Estimation. From the perspective of hardware-specific resource analysis, QFTs are likely to feature in application-relevant resource analysis, yet can be scaled down to a small handful of logical input gates, allowing us to trace and debug compilation workflows in our tool chain at the level of logical operations. This makes QFT circuits ideally suited as first test cases of our estimation framework. This work analyzes QFT circuits on $n$ qubits where $n=4,10,25,50,75,100,250,500$ or $1000$.
II.2 Hamiltonian simulation
Simulation of lattice-based Hamiltonians covers a family of critical applications investigated by researchers in quantum chemistry and condensed matter. Moreover, simulating dynamics may provide utility for smaller hardware system sizes than methods for evaluating static properties, such as the ground state.
Hamiltonian simulation consists of evolving an initial state |$\ket{\psi_{0}}$ to a state at time $t$, |$\ket{\psi}$ according to,
|$\ket{\psi}=e^{-iHt}\ket{\psi_{0}}$ (23)
where $H$ is the Hamiltonian of the system of interest.
We consider the simulation of two families of Hamiltonians: 2D Transverse-Ising models and 2D Fermi-Hubbard models. Each of these will be described in more detail below.
II.2.1 Transverse-Ising Hamiltonian Simulation
The Los Alamos National Laboratory (LANL) has created a public repository containing circuit generation and analysis of problems that may be amenable to quantum computing, and ostensibly are of interest to the scientific community, [72]. Drawing inspiration from their investigation of magnetic lattices, we perform resource estimation for the simulation of a time-invariant Transverse-Ising Hamiltonian,
$H=\sum_{\langle\textbf{jk}\rangle}\,Z_{\textbf{j}}\otimes Z_{\textbf{k}}+0.1\sum_{\textbf{j}}\,X_{\textbf{j}},$ (24)
on a triangular lattice, shown in Figure 8.
Figure 8: Triangular lattice of size $4\times 4$
Here, $X_{\textbf{i}}$ and $Z_{\textbf{i}}$ represent the Pauli X and Z operators, respectively, acting on lattice site i, and angle brackets denote lattice sites connected by an edge.
Based on the studies documented by LANL, we carried out resource estimates of simulation of the Hamiltonian in eq. 24 for a total time of $t=1$ (in the inverse Hamiltonian units) on a 2D triangular lattice of size $N\times N$ where $N$ is an integer between $2$ and $19$.
II.2.2 Fermi-Hubbard Hamiltonian simulations
We also investigated the Fermi-Hubbard simulation as a core computational capability for solving room-temperature superconductivity problems. These problems typically involve studying the ground state energy, state configurations, and other observable system properties, including time-dependent quantities.
The Fermi-Hubbard Hamiltonian we consider is given by,
$H=$ $-J\sum_{\langle\textbf{j},\textbf{k}\rangle,\sigma}\left[c^{\dagger}_{\textbf{j},\sigma}c_{\textbf{k},\sigma}+c^{\dagger}_{\textbf{k},\sigma}c_{\textbf{j},\sigma}\right]$
$+J^{\prime}\sum_{\langle\langle\textbf{j},\textbf{k}\rangle\rangle,\sigma}\left[c^{\dagger}_{\textbf{j},\sigma}c_{\textbf{k},\sigma}+c^{\dagger}_{\textbf{k},\sigma}c_{\textbf{j},\sigma}\right]$
$+U\sum_{\textbf{j}}n_{\textbf{j},\uparrow}n_{\textbf{j},\downarrow}$
$+\frac{h_{z}}{2}\sum_{\textbf{j}}(n_{\textbf{j},\uparrow}-n_{\textbf{j},\downarrow})$
$+\mu\sum_{\textbf{j},\sigma}n_{\textbf{j},\sigma},$ (25)
where single and double angled brackets denote all nearest neighbor and next-nearest-neighbor lattice site pairs respectively. $c^{\dagger}_{\textbf{j},\sigma}(c_{\textbf{j},\sigma}$) creation (annihilation) operators of a spin-$\sigma$ Fermion on the jth lattice site where $\sigma\in\{\uparrow,\downarrow\}$. $J$ (primed or unprimed) and $U$ capture the magnitudes of the hopping and repulsion energies respectively. $n_{\textbf{k},\sigma}=c^{\dagger}_{\textbf{k},\sigma}c_{\textbf{k},\sigma}$ is the number operator, $h_{z}$ is an external applied magnetic field, and $\mu$ is the chemical potential.
To motivate the specific Fermi-Hubbard simulations, we will obtain resource estimates. We will consult the literature to understand configurations that interest the community. Broadly speaking, methods for simulating a Fermi-Hubbard system can be broken into stochastic and deterministic families. As mentioned in [73], deterministic methods do not “suffer from numerical issues seen in stochastic simulations when dealing with systems with non-zero chemical potential”. Moreover, this instability due to the “sign problem” becomes prohibitive for moderate $\mu\sim 0.3$. This suggests separating instances into those with $\mu=0$, and those with $\mu>0.3$. Here, we reference a few exemplary studies in the literature on stochastic and deterministic methods.
[…]
Of note is that in [73] it was determined that the time required to run these studies on a CPU varies like,
$t_{\text{C}PU}\propto L_{x}L_{y}(2\chi^{3}D^{4}d^{2}+\chi^{2}D^{2}d^{2}+\chi^{2}D^{4}d^{4}),$
where $\chi$ is the truncated bond dimension and $d$ the problem dimension. Memory scaled as,
$2\chi^{2}D^{4}d^{2}M_{\text{n}um}+2N\chi^{2}D^{2}dM_{\text{n}um}.$
The authors of the study empirically found $\chi=3D$ was sufficiently accurate for their studies, with error about $1\%$. These estimates provide an interesting guidepost for the tipping point at which FTQCs quantum computers may outperform classical computers for these types of problems.
Taking these studies as examples of the types of systems of interest to the community it suggests obtaining resource estimates for Fermi-Hubbard Hamiltonian simulations consisting of a 2D square lattice for total time $t=1$, with $J=1$, $J^{\prime}=0$, $h_{z}=0$, $\mu=0$, $U=2$, and lattice sides of length $N=2$ to $20$, $30,50,75$ and $100$.
II.2.3 Algorithm and Approach
Owing to its optimal scaling characteristics, in terms of queries to the block encoding of the Hamiltonian [14], we have chosen Quantum Signal Processing (QSP) as the Hamiltonian simulation algorithm for which we obtain a resource estimate. QSP is briefly described below, but more details can be found in [14], [78], [79], and [80].
Given a Hamiltonian $H$ and simulation time $t$, QSP performs a sequence of single qubit rotations, parameterized by a sequence of phase angles, interlaced with a quantum walk-like operator. This allows one to apply a polynomial approximation of $e^{-iHt}$ to the input target state. The sequence of phase angles is responsible for the polynomial approximation. It is computed classically offline, according to the error tolerance of the approximation error desired, captured by the user-defined parameter $\epsilon_{\text{qsp}}$. $\epsilon_{\text{qsp}}$ is also inversely related to the probability of success of the QSP algorithm. Thus, for an accurate simulation, with a high probability of success, $\epsilon_{\text{qsp}}$ may have to be set relatively small.
For implementation details, we follow the examples accompanying the pyLIQTR [81] package (v1.3.3), as well as the work done in [82, 83].
For demonstration purposes, we set the desired probability of success of our algorithm, $p_{\rm algo}$, to be equal to 0.95. In real world applications this is typically fixed to meet outcome requirements (such as accuracy, wall clock time, etc). Denoting the probability of success of a QSP circuit, $p_{\rm qsp}$, we therfore require,
$p_{\rm algo}\leq p_{\rm qsp}.$ (26)
Since circuit size is inversely related to $\epsilon_{\text{qsp}}$, it is prudent to choose $\epsilon_{\text{qsp}}$ to be as large as possible while still guaranteeing that $p_{\rm qsp}$ exceeds the algorithm target of 0.95. According to [14] $(1-2\epsilon_{\text{qsp}})\leq p_{\rm qsp}$. Thus, to ensure eq. 26 is satisfied we set,
$1-2\epsilon_{\text{qsp}}=p_{\rm algo},$ (27)
and so in this work we choose $\epsilon_{\text{qsp}}$ to be equal to 0.025.
II.3 Circuit Synthesis
For our Hamiltonian simulation circuits, we used pyLIQTR v1.3.3 to construct the Hamiltonians and synthesize the QSP circuits. The pyLIQTR package provides tools that construct the Hamiltonians, calculate the necessary phase angles for QSP simulation, and generate QSP circuits. In addition, pyLIQTR allows the user to request random phase angles rather than solving for the precise values required for an accurate simulation. In our work, we used random phase angles. This is appropriate because we only perform resource estimation, not numerically accurate simulations. The correct number of phase angles is used in the circuits, but their numerical values are random. This provides a conservative overestimate of the resource estimations without explicitly incurring the time-consuming step of solving for the phase angles.
For Fermi-Hubbard models, we utilized pyLIQTR’s FermiHubbard model class, and employed the FermiHubbardSquare block encoding, which is based on [84].
The same overall approach was taken for the time evolution of the Transverse-Ising models. However, for these problems, we used pyLIQTR’s Heisenberg model class, and PauliLCU block encodings. For QFT instances, we used the QFT circuits provided by the qft method in cirq.
III Resource estimation results
Key physical resource requirements for the test cases described in section II, over various problem sizes, are summarized in table 2. We demonstrate that our tool can estimate resources for problem sizes consistent with studies in the literature and beyond the scale of current studies.
Our methodology allows us to extend the state of the art in a few ways. (1) A full accounting for resources for both T counts, but also for executing the Clifford portions of the circuit as well. (2) A breakdown of physical and temporal resources allocated to portions of the execution of the algorithm, namely: Clifford, T-state and teleportation between modules. (3) A quantitative estimation of the impact on total resources of the performance of subsystems.
When we visualize this ensemble of compiled logical resources and physical resource requirements (fig. 9 and fig. 10 respectively), we see the run time and ancilla bus size trajectories that show an increasing trend in resources with increasing problem size. This provides us with some validation of the reasonableness of our methodology and its benefit as a tool to support studies into the future. Interestingly, we show a decreasing trend in logical qubits for small-sized problems until the number of modules increases. Once the number of modules increases, we see a sharp increase in the number of logical qubits. This is because physical qubits are assigned to T factories instead of logical qubits or the ancilla bus. As the problem size grows, the number of qubits used in the bilinear quantum memory and ancilla bus pattern increases in the single-module regime. This reduces the number of T factories that can fit in the module. As the T factories are large, monolithic regions of logical qubits, removing them gives a net reduction in the number of logical qubits in use, even though the number of patches in the bilinear pattern is larger. The storage space occupied in a room can decrease even as smaller objects are added, provided those smaller objects require the removal of large objects to accommodate them.
Table 2: Rigetti’s resource estimation results on a superconducting fault-tolerant quantum computing architecture with space-optimized graph-state compilation. Here, the Logical Bus includes patches for the memory and the ancilla bus mediating logical measurements in the bilinear pattern of fig. 6. Likewise, $T$ Patches includes logical patches for all $T$ Distillation Factories and $T$ Buffers.
Input circuit Logical Resources Superconducting FTQC resources
Instance logical qubits Logical Bus T Patches Physical qubits Modules $t_{\text{tot}}^{\text{hardware}}$ $E_{\text{tot}}$
Fermi Hubbard
2x2 37 276 4.05k 1.7M 2 4.16s 679Wh
4x4 74 332 3.22k 1.82M 2 37.5s 6.12kWh
8x8 182 728 2.07k 1.81M 2 9.5m 93kWh
16x16 578 2.31k 1.25k 3.45M 4 6.9h 8.11MWh
20x20 878 3.52k 1.4k 5.21M 6 37.8h 66.7MWh
50x50 5.05k 20.2k 5.54k 37.6M 42 7.13y 772GWh
100x100 10.1k 40.7k 24k 141M 174 85.9ky 38.5TWh
QFT
N-4 4 88 6.38k 1.56M 2 2.8ms 457 mWH
N-10 10 248 5.5k 1.66M 2 41.7ms 6.81Wh
N-25 25 572 3.75k 1.7M 2 427ms 69.7Wh
N-100 100 3.39k 2.76k 3.56M 4 10.4s 3.39kWh
N-500 500 18.4k 4.5k 18.3M 20 2.03m 199kWh
N-1000 1k 37.2k 9.57k 45.3M 52 10.6m 2.69MWh
Transverse-Ising
2x2 67 636 2.43k 1.77M 2 64.4s 10.5kWh
5x5 212 2.42k 816 3.43M 4 6.75h 7.92MWh
10x10 654 10.1k 2.78k 17.4M 20 5.07d 715MWh
19x19 2.22k 41.1k 13.5k 92M 110 82.6d 64.1GWh
Figure 9: Fault-tolerant compiled logical qubits and T count for three sets of test cases. Grey indicates Quantum Fourier Transform results for sizes 4 to 1000. The black plot is the Hamiltonian evolution of the Transverse-Ising model on triangular lattice sizes ranging from 2x2 to 19x19. The teal line is estimates for Hamiltonian simulation of a Fermi-Hubbard model on a square lattice with sizes ranging from 2x2 to 100x100.
Figure 10: Fault-tolerant hardware runtime and required number of physical qubits in the bilinear pattern for three sets of test cases. Grey indicates Quantum Fourier Transform results for sizes 4 to 1000. The black plot is the Hamiltonian evolution of the Transverse-Ising model on triangular lattice sizes ranging from 2x2 to 19x19. The teal line is an estimate for the Hamiltonian simulation of a Fermi-Hubbard model on a square lattice with sizes ranging from 2x2 to 100x100. We set the Hamiltonian evolution time to 1 in units of $1/J$ for both the Transverse-Ising and Fermi-Hubbard models.
(a) QFT
(b) Transverse-Ising
(c) Fermi-Hubbard
Figure 11: Physical qubit allocations for the types of qubits described in the microarchitecture shown in fig. 1; those involved in T state distillation, logical register and ancillas for queuing and routing.
(a) QFT
(b) Transverse-Ising
(c) Fermi-Hubbard
Figure 12: Proportion of quantum computer compute time allocated to graph state consumption (including T state injection), intermodule communications, and decoding delay for time evolutions.
Figure 13: Comparison of normalized fault-tolerant runtime and number of intermodule pipes. The number of intermodule pipes varies from 1 to 32. Run time is normalized to the value with one pipe. Black indicates Quantum Fourier Transform results of size 10. The grey plot is the Hamiltonian evolution of the Transverse-Ising model on a triangular lattice of size 10. The teal line captures estimates for Hamiltonian simulation of a Fermi-Hubbard model on a square lattice with size 10x10.
Figure 14: Total physical qubits required for Hamiltonian evolution of the Fermi-Hubbard of various sizes. Line colors denote different decoders; Teal indicates Minimum Weight Perfect Matching (MWPM), black is belief propagation, grey is belief propagation with ordered statistics decoding, and magenta is Astra [85]
The full accounting of both Clifford and T contributions to total resources, (1), is shown in fig. 11 for all families of applications for which we estimated resources. It is interesting to note that conventional T-counting methodologies don’t often surface the competition for resources between the Clifford and T-state generation of a quantum architecture. As problems scale, physical qubits need to be allocated to the increased needs of Clifford processing, squeezing out available qubits that could be allocated to MSD. This, in turn, reduces the number of MSD factories that can be applied to a calculation and thus increases runtime in a non-linear fashion.
This impact on the breakdown of total runtime, (2), can be seen in fig. 12. Here, the contribution to runtime of graph consumption (including T state injection and graph state consumption) increases dramatically for the largest applications considered. This is because fewer qubits are allocated to create higher precision T-states, doubling the penalty in runtime. The impact of runtime on the teleportation present in the distributed architecture can also be seen in this figure. Despite a relatively slow inter-module interaction time (5x slower than the native code cycle), the contribution to overall runtime is application dependent. For QFT and transverse Ising applications, a relatively small portion of the overall runtime is spent performing this intermodule interaction, even at large problem sizes. This shows that computation can be partitioned for specific architectures and algorithms to allow distributed computing, with relatively low impact. However, this seems algorithm-dependent, and for Fermi-Hubbard problems, the penalty paid for distributed quantum computation is high.
Moreover, we can test the sensitivity of runtime to parameters in this architectural model, (3). In fig. 13 we select an instance from each family, vary the number of inter-module connections, and calculate the impact on runtime. We define this in terms of a pipe where the number of connections equals the required code distance for the algorithm. Additional pipes will allow the teleportation of nodes between widgets to happen in parallel. This is done to quantify the penalty being paid for not having a fully connected lattice of qubits between modules. Here we see that while there is indeed an increased runtime for a reduced number of inter-module connections, there are diminishing returns for adding additional connections. Once again, this impact is seen to be highly application-dependent.
Similarly, we can directly quantify the impact of changing the decoders on the physical system size. Results for Fermi-Hubbard instances of increasing problem size are shown in fig. 14 and reflect the utility and flexibility of our resource estimation tool to asses the impacts of changing multiple subsystems within the overall architecture.
IV Future Work
The results from Section III can be improved by focusing on at least the following two aspects: a) using a decoder which is faster and/or has better scaling than the standard choice of MPWM; b) optimizing the widget circuits by, for example, reducing circuit gate counts and depths. The savings potential of these two approaches will be detailed in the following.
IV.1 Replacing the MWPM decoder
The MWPM decoder (e.g.,  [86]) is the de facto decoder for performing resource estimates. Other decoders have been proposed, but MWPM is still the most efficient regarding the decoding speed vs. physical error rate trade-off: MWPM can handle high error rates while its complexity is polynomial. There are faster, lower polynomial complexity decoders (e.g., union-find variations or pure belief-propagation) and decoders with linear complexity, which are challenging to scale to high distances (machine learning decoders). Nevertheless, to the best of our knowledge, no alternative decoder exists that has, at the same time, a lower degree of polynomial decoding complexity and a threshold higher than MWPM.
Herein, we report results, fig. 15, about a machine learning decoder that outperforms MWPM: it has a higher threshold and linear runtime complexity. Our machine learning decoder, Astra [85], uses graph neural networks (GNNs) in a novel way: we are learning a message-passing algorithm on the Tanner graph of the codes. We can train, within hours, such decoders with gaming GPUs for surface codes up to distance 11. In contrast to other machine learning decoders, the training time of our decoder is significantly lower (e.g. [87]) for comparable depths.
Figure 15: The Logical Error Rate (LER) for code capacity depolarising noise of Astra vs MWPM. Astra has a threshold of $\sim 17\%$, and MWPM has a threshold of $\sim 14\%$. Astra clearly outperforms MWPM in terms of LER. In fact, Astra d9’s results are better than MWPM d11’s.
We train distinct GNN decoders for each distance using syndromes collected by passing the surface codes through a code capacity noise channel. This channel assumes that the data qubits are affected by noise with a probability $p$ and that the syndrome measurements are perfect. For this work, we do not report the decoding of circuit-level correlated noise (e.g., [86]) and focus entirely on correcting depolarizing noise.
After training our decoders, we evaluate their decoding performance numerically and capture the decoder’s performance through an equation of the type from Eq. 1:
$p[{\mathrm{GNN}}]=0.56\left(\frac{p}{0.17}\right)^{\frac{d+1}{2}}~{}.$ (28)
Usually, the performance scaling equations for the surface code are obtained by simulations with circuit level noise (e.g., eq. 1) because this is considered the most exact model. Nevertheless, the code capacity is more straightforward for benchmarking the preliminary decoder versions. To allow the comparison between MWPM and Astra, we also benchmark MWPM with code capacity noise. The results are illustrated in Fig. 15. Based on the MWPM code capacity errors, we obtain the MWPM scaling equation as
$p[{\mathrm{MWPM}}]=0.52\left(\frac{p}{0.14}\right)^{\frac{d+1}{2}}~{}.$ (29)
We observe that Astra has a higher threshold compared to MWPM ($17\%>14\%$) and offers faster (better) logical error suppression ($0.56>0.52$). This implies that, by replacing the MWPM with Astra, one can achieve the same logical error rate with lower distance surface codes. After benchmarking its performance under circuit-level noise, future work will focus on quantifying the exact savings when utilizing the new decoder.
IV.2 Ultra-large scale circuit optimization
The Fermi-Hubbard circuit instances analyzed in the previous sections are, for practical regimes, too large in terms of gate counts. Compiling, analyzing, and optimizing such circuits at graph-state or Clifford+T level is challenging: the state-of-the-art quantum software tools cannot handle ultra-large-scale circuits.
Widgets (Section A.2.4) are subcircuits representing relevant parts of the larger circuits. Due to their size, widgets can be more easily analyzed and optimized. Nevertheless, even widgets are complicated to optimize for metrics such as $T$-count and $T$-depth. Optimization is a highly complex task, and most optimization heuristics are not designed (or do not perform as expected) for (sub)circuits of tens of qubits and hundreds of gates.
Tools like pyLQTR are built on top of other quantum software frameworks, such as Google Cirq, which were not designed for extreme scalability. We bridge this gap by developing and implementing a software tool that can easily handle ultra-large-scale circuits. Our tool, Pandora  [^1] https://github.com/ioanamoflic/pandora/, is based on PostgreSQL and leverages the capabilities of modern relational database software supporting multi-threaded operations on quantum circuits.
Pandora offers a simple API that can be used to develop new and more complex optimization methods. We interfaced our tool with Google Qualtran [89] and can easily extract large circuits for analysis and optimization. Currently, the circuits and widgets available in Qualtran have been optimized manually, so automatic optimization is even more challenging. Therefore, the results from Table 3 belong to far from optimal circuits.
Table 3 reports preliminary results on optimizing arithmetic widgets as presented in [90]. The latter are ubiquitous building blocks of practical application circuits like the Fermi-Hubbard. Our optimization procedure for optimizing the circuits is not as complex as the one described by Ref. [90]. We implemented a heuristic that uses multiple threads for applying circuit template rewrites [90]: 1) decomposition of gates (e.g., Toffoli); 2) canceling H, T, X gates; 3) canceling CNOT gates; 4) commuting gates to the left and the right; 5) reversing directions of CNOT gates.
To conclude, quantum software for ultra-large-scale circuit analysis and optimization can potentially lower the resources required for practical quantum computing applications. Pandora is a first step towards optimizing at scale.
Circuit Orig. T-count Opt. T-count Impr.
Adder8 266 236 11.27%
Adder16 602 536 10.96%
Adder32 1274 1144 10.20%
Adder64 2618 2368 9.54%
Adder128 5306 4648 11.64%
Adder256 10682 9258 13.33%
Adder512 21434 18562 13.39%
Adder1024 42938 38216 10.99%
Adder2048 85946 76056 11.50%
Table 3: Widget optimization results: addition circuits from Maslov [90] are optimized for T-count. The original T-count is reported in the second column, and the optimized T-count in the third column. The optimization of Adder2048 required approx. one hour.
V Conclusion
This paper proposes a fault-tolerant architecture based on modular quantum processing units utilizing superconducting qubits and the graph state formalism. We provide quantitative estimates for the scale and runtime of executing various quantum algorithms at the utility scale, employing a resource estimation framework and software tool. This framework estimates the physical resources required for quantum algorithms and offers estimates for test cases derived from existing literature. We have demonstrated that a fault-tolerant quantum computer built on a distributed superconducting architecture, consisting of two legs of modules, could support scientifically interesting applications within a reasonable runtime. However, a significant algorithm-dependent penalty exists for distributing computation across multiple modules. These estimates align with established results, such as factoring [2, 6]. This model-driven approach delineates how a superconducting platform could accomplish this, including details on scheduled execution over the preparation and consumption of graph states and hardware-level justifications for some of the most critical size and power considerations.
This work has also enabled research toward ensuring that real quantum computing platforms can be developed to provide the advantages of fault-tolerant operation. It serves as a software testbed for evaluating the impact of algorithmic, compilation, decoding, and physical-layer intrinsic or system architecture proposals. Moreover, this allows for direct per-algorithm comparison of different modalities by assessing run times, spatial, and energy resource costs. Rapid iteration in a simulated environment is a proven method for accelerating technology growth, and we hope that our efforts to produce this tooling contribute positively to this endeavor.
VI Acknowledgements
The views, opinions, and/or findings expressed are those of the author(s). They should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. This research was developed with funding from the Defense Advanced Research Projects Agency under Agreement HR00112230006. Jannis Ruh was supported by the Sydney Quantum Academy, Sydney, NSW, Australia.
Appendix A Supplemental Material
A.1 Glossary of default RRE outputs
With the workflow of the graph-state-based fault-tolerant resource estimation method outlined in the main text, we summarize and reiterate the selected output parameters of RRE, providing new details in some cases. The parameters listed are RRE outputs to stdout and CSV files by default. We detail 48 logical architecture or hardware-relevant output parameters and explain how to calculate them. The delivered CSV file can include results for additional output parameters. Users can add custom classes of quantities for estimation to RRE using class templates from the resources.py module.
  • Parameter 1: Code distance, $d$
    $d$ denotes the surface code distance. All logical patches have a size of $2d^{2}$ and serve as foundational elements for all architectural components in the FTQC.
    We need to calculate the minimum viable $d$ precisely and iteratively according to Algorithm 1, which requires satisfying a logarithmic inequality in Step 6, also referred to in eq. 17.
  • Parameter 2: Number of logical qubits in the quantum memory register, $\tilde{n}_{\text{logical}}$
    The term $\tilde{n}_{\text{logical}}$ refers to the maximum number of logical nodes needed to process the complete graph state $\mathcal{G}$ according to the widgetization-based set $\tilde{S}_{\text{est}}$ at any given time. In other words, this represents the maximum quantum memory required. Each complete quantum memory register chain, shared among the modules on the same leg of the macro-architecture, contains $\tilde{n}_{\text{logical}}$ logical qubits. To prevent any quantum memory bottlenecks or delays, we allocate a fixed quantum memory and ancilla bus space of size $2\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$ to all modules and steps of the schedules. There may be more efficient methods for assigning the bilinear quantum register and ancilla bus spaces.
    Note that the maximum quantum memory required for the complete $U$ is a distinct quantity, $n_{\text{logical}}$. For widgetized circuit input, we derive a proxy for this quantity from the subgraphs as $\tilde{n}_{\text{logical}}=\max(n^{g_{0}}_{\text{logical}},\cdots,n^{g_{n_{\text{widgets}}-1}}_{\text{logical}})\approx n_{\text{logical}}$. Cabaliser can efficiently output each $n^{g}_{\text{logical}}$ for sufficiently small graph states.
  • Parameter 3: Number of $T$-distilleries (MSD factories) per module, $n_{\text{distilleries}}$
    The number of Litinski’s $T$-state distillation widgets or MSD factories per module is calculated as $n_{\text{distilleries}}=\lfloor\frac{l_{\text{edge}}-1}{\lfloor\frac{l_{\text{width}}}{2d^{2}}\rfloor}\rfloor n^{\text{col}}_{\text{distilleries}}$. The factories send purified magic states to $T$-transfer and ancillary buses necessary for the measurements on the logical qubits during the consumption stage. For more details, see section I.4.6.
  • Parameter 4: Number of logical memory qubits per module
    The number of logical memory, or graph-state, or data qubits in the bilinear pattern in every module is equal to $\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$. These correspond to the blue tiles in fig. 6. In other words, every module on a leg of the macro-architecture contains this same number of logical memory qubits. Although the last module may require fewer active nodes for quantum operations, the same space is always allocated to this component.
  • Parameter 5: Number of physical memory qubits per module
    The number of memory, or graph-state, or data physical qubits in the bilinear pattern in every module is equal to $2d^{2}\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$.
  • Parameter 6: Number of logical ancilla qubits in the ancilla bus per module
    The number of logical ancilla qubits in the ancilla bus section of each module is equal to $\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$. These correspond to the yellow tiles in fig. 6, which FTQC uses to facilitate logical measurements on the graph-state qubits. Each module on a leg of the macro-architecture contains the same number of logical ancilla qubits. Although the last module may need fewer active nodes for quantum operations, the same space is always allocated for this component.
  • Parameter 7: Number of physical ancilla qubits in the ancilla bus per module
    The number of physical ancilla qubits in the ancilla bus portion of every module is equal to $2d^{2}\lceil\tilde{n}_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil$.
  • Parameter 8: Number of logical qubits in the $T$-transfer bus per module, $l_{\text{transfer-bus}}$
    The number of logical qubits in the $T$-transfer bus portion of every module is calculated as $l_{\text{transfer-bus}}=(l_{\text{edge}}-l_{\text{qbus}}-n^{\text{col}}_{\text{distilleries}}\lceil\frac{L_{\text{length}}}{\sqrt{2}d}\rceil)l_{\text{edge}}+n^{\text{col}}_{\text{distilleries}}\lceil\frac{L_{\text{Length}}}{\sqrt{2}d}\rceil$. These represent the teal tiles in fig. 6, which FTQC utilizes to actively store and transfer purified $T$-states from MSD factories to the ancilla bus.
  • Parameter 9: Number of physical qubits in the $T$-transfer bus per module
    The number of physical qubits in each module’s $T$-transfer bus portion is $2d^{2}l_{\text{transfer-bus}}$.
  • Parameter 10: Number of logical qubits allocated to all $T$-distillation factories in each module
    The number of logical nodes assigned to all $n_{\text{distilleries}}$ factories in each module can be calculated as $n_{\text{distilleries}}\lceil L_{\text{length}}/(\sqrt{2}d)\rceil\lceil L_{\text{width}}/(\sqrt{2}d)\rceil$. These correspond to purple tiles in fig. 6.
  • Parameter 11: Number of physical qubits allocated to all $T$-distillation factories in each module
    The number of active physical qubits across all $n_{\text{distilleries}}$ factories per module can be computed as $n_{\text{distilleries}}Q$. After choosing a factory from the Distillation Widget Look-up table 1, $Q$ becomes the number of physical qubits the factory requires for MSD.
  • Parameter 12: Total number of modules in both legs of macro-architecture
    The total number of interconnected QPU cryo-modules on both legs of the macro-architecture can be computed as $2n^{\text{per-leg}}_{\text{modules}}$. The parameter $n^{\text{per-leg}}_{\text{modules}}$ is the number of modules per leg or the height of the ladder structure. See fig. 3 for details.
  • Parameter 13: Total number of interconnects in the macro-architecture
    The total number of width-$d$ coherent interconnects in the ladder macro-architecture can be computed as $n_{\text{inter-pipes}}(3n^{\text{per-leg}}_{\text{modules}}-2)$. FTQC uses these connections to perform inter-module graph preparation operations on the same-leg or teleporting output to input nodes on the opposite-leg modules. See fig. 3 for details.
  • Parameter 14: Number of allocated physical qubits per module
    The number of allocated physical qubits per module for all components can be computed as $2\lceil n_{\text{logical}}/n^{\text{per-leg}}_{\text{modules}}\rceil d^{2}+n_{\text{distilleries}}Q+2l_{\text{T-transfer}}d^{2}$.
  • Parameter 15: Number of algorithmic logical qubits, $n_{\text{input}}$
    The number of algorithmic logical qubits or width, $n_{\text{input}}$, of the complete input circuit, $U$. For the time-sliced widgetization we perform on the input algorithm, cf. fig. 7, all widgets have the same number of output and input nodes equal to $n_{\text{input}}$.
  • Parameter 16: Diamond-norm precision for gate synthesis, $\varepsilon$
    Extensive analytical and numerical studies (see, for example, [68, 69]) have established that a non-Clifford 1Q rotation gate can be systematically approximated as a chain of Clifford+$T$, a process known as gate synthesis. Here, the upper bound for $T$-length scales as $L_{\varepsilon}=\lceil c_{0}\log_{2}(1/\varepsilon)+c_{1}\rceil$. The diamond-norm precision, $\varepsilon$, indicates the rate at which FTQC utilizes the gate synthesis decomposition of arbitrary-angle rotations to perform $Rz$-basis measurements at the end, during each consumption step. We assume the use of the state-of-the-art decomposition approach of mixed-fallback [69] by default, which offers $c_{0}=0.57,c_{1}=8.83$, representing an improvement over the previously employed “GridSynth” method [68], which had $c_{0}=3.0,c_{1}=0$.
    RRE processes the arbitrary angle decompositions internally, which means the constructed graphs do not depend on $\varepsilon$-value and can be cached. The diamond-norm precision must be determined accurately and iteratively, following Algorithm 1, which naturally involves solving a logarithmic equation in Step 6.
    We argue that $\varepsilon$ should not be left for the user to set arbitrarily (although RRE allows users to impose a fixed value if needed). Extreme caution is required here because precisions are, indeed, equivalent to the fault-tolerant resources required: if $\varepsilon$ is chosen too large, the rate condition in Step 6 of Algorithm 1 is not met, while if it’s too small, the space-time resources will become wasteful and excessively large.
  • Parameter 17: Total number of T-gates, or T-count, or $N^{tot}_{T}$
    In the graph-state-based estimation method, our proxy for the total T-count, or $N^{\text{tot}}_{T}$, is the total number of T-basis measurements required after the arbitrary-angle $R_{z}$’s are synthesized at the measurement points of the consumption stage. Therefore, we can write T-count as $N^{\text{tot}}_{T}=N^{Rz}_{\text{init}}L_{\varepsilon}+N^{T}_{\text{init}}=N^{Rz}_{\text{init}}\lceil c_{0}\log_{2}(1/\varepsilon)+c_{1}\rceil+N^{T}_{\text{init}}$.
  • Parameter 18: Effective $T$-depth
    This parameter estimates the effective $T$-depth for the input logical circuit, $U$. This can be computed as $\lceil N^{T}_{tot}/\tilde{n}_{\text{logical}}\rceil$.
  • Parameter 19: Number of $Rz$-gates in the input circuit, $n^{\text{init}}_{Rz}$
    The number of arbitrary-rotation $Rz$-gates in the complete input algorithm, $U$, whether written explicitly or contained within other logical gates. All $Rz$ gates must undergo gate-synthesis decomposition [68, 69] to Clifford+$T$ at the end, during the consumption stage.
  • Parameter 20: Number of $T$,$T^{\dagger}$-gates in the input circuit, $N^{\text{init}}_{T}$
    The number of $T$ and $T^{\dagger}$ gates in the complete input algorithm, $U$, whether written explicitly or contained within other logical gates. We perform this count before performing gate-synthesis Clifford+$T$ decomposition of arbitrary-angle $R_{z}$-gates.
  • Parameter 21: Number of Cliffords in the input circuit, $N^{\text{init}}_{\text{Clifford}}$
    The number of explicit Clifford gates in the complete input algorithm, $U$.
  • Parameter 22: Number of graph state nodes, $N$
    $N$ represents the total number of logical nodes in the complete graph state, $G$, or the graph size. For widgetized circuits, we calculate a proxy for this quantity from subgraphs as $\tilde{N}=\sum_{i=0,...,n_{\text{widgets}}-1}N_{i}+\sum_{i=0,...,n_{\text{widgets}}-2}n^{\text{outputs}}_{i}\approx N$. Here, $n^{\text{outputs}}_{i}=n_{\text{input}}$ denotes the number of output nodes at step $i$ that must be added to the graph stitched from subgraphs to enable Bell pair teleportations.
  • Parameter 23: Total number of measurement steps in the consumption schedule
    The total number of sequential measurement steps in the consumption schedule over all widgets (repeated or not). In other words, the total number of subsets in the nested sets in the consumption schedule outputted by Cabaliser[71]. This quantity governs the quantum operation runtime for the graph consumption stage.
  • Parameter 24: Total number of measurement steps in the preparation schedule, $\mathbb{L}^{\text{prep}}$
    The total number of sequential measurement steps in the preparation schedule over all widgets (repeated or not). In other words, the total number of sub-sets in the nested sets in the preparation schedule outputted by Substrate Scheduler[44]. This quantity governs the quantum operation runtime for the graph preparation stage.
  • Parameter 25: Total number of widgets or widgetization time steps, $n_{\text{widgets}}$
    The total number of widgets (repeated or not), which is equivalent to the number of time steps in $U$ given the time-sliced decomposition presented in fig. 7 and section A.2.4.
  • Parameter 26: Number of distinct widgets, $n^{*}_{\text{widgets}}$
    $n^{*}_{\text{widgets}}$ is the number of distinct widgets in set $[U_{0},\dots,U_{n_{\text{widgets}}-1}]$ for the time-sliced widgetization of $U$. This means one can always write $n^{*}_{\text{widgets}}\leq n_{\text{widgets}}$. This is the number of widgets that FTQC needs to compile.
  • Parameter 27: decoder tock, $t^{\rm decoder}_{\rm tock}$
    For our proposed superconducting fault-tolerant architecture, we assume it is feasible to engineer generic decoders utilizing modern GPU or FPGA hardware and methods such as MWPM or neural networks. $t^{\rm decoder}_{\rm tock}$ is the time required for the decoder to complete decoding $d$ QEC sweeps expressed in seconds. This can be approximated as $t^{\rm decoder}_{\rm tock}=dt_{\text{decoder}}$. We assume it is feasible to engineer generic GPU or FPGA-based decoders with a decoding tock or delay in the same order of magnitude as $t^{\rm decoder}_{\rm tock}$ for superconducting architecture. Here, we consider a characteristic decoder cycle time of $t_{\text{decoder}}=1~{}\mu$s.
  • Parameter 28: Intra-module quantum tock for graph processing, $t^{\rm intra}_{\rm tock}$
    The time needed for FTQC to sequentially perform $d$ QEC sweeps for all intra-module quantum operations, which prepare or consume the subgraphs, is computed as $t_{\rm intra}=8dt$. Here, $t$ represents the characteristic (dominant) gate time that sets the architectural tock for all intra-modular operations. The factor of 8 arises from the fact that in our physical architecture, each surface code sweep for syndrome extraction circuits necessitates one initialization, two Hadamards, four entangling gates, and one measurement, all of which are assumed to take a duration of $t$ (see, for example, Fig. 9 of [10]).
  • Parameter 29: Intra-module quantum tock for $T$-state distillations
    The time required for MSD factories to distill a $T$ state performing $d$ QEC sweeps sequentially. This can be computed as $8Ct$ for factories built from rotated surface code patches.
  • Parameter 30: Total number of available physical qubits, $n_{\rm avail-phys}$
    Total number of physical qubits available in all $2n^{\text{per-leg}}_{\text{modules}}$ modules of the FTQC.
  • Parameter 31: Number of available logical qubits per module, $n_{\text{avail-logical}}$
    The number of logical qubits available per module allocated to different architectural components and quantum operations. This can be computed as $n_{\text{avail-logical}}=l_{\text{edge}}^{2}$.
  • Parameter 32: Total number of unallocated logical qubits
    Considering all modules and operations of the FTQC, the total number of unallocated logical qubits. This parameter can be computed as $2n_{\text{unalloc-logical}}n^{\text{per-leg}}_{\text{modules}}$, where $n_{\text{unalloc-logical}}=l_{\text{edge}}^{2}-2\lceil\frac{\tilde{n}_{\text{logical}}}{n^{\text{per-leg}}_{\text{modules}}}\rceil-l_{\text{T-transfer}}-n_{\text{distilleries}}\lceil\frac{l_{\text{length}}}{\sqrt{2}d}\rceil\lceil\frac{l_{\text{width}}}{\sqrt{2}d}\rceil$.
  • Parameter 33: Total number of unallocated physical qubits
    Considering all modules and operations of the FTQC, the total number of unallocated physical qubits. This parameter can be computed as $4n_{\text{unalloc-logical}}n^{\text{per-leg}}_{\text{modules}}d^{2}$.
  • Parameter 34: Total number of allocated logical qubits
    Considering all modules and operations of the FTQC, the total number of allocated logical qubits to different architectural components. This parameter can be computed as $2n_{\text{alloc-logical}}n^{\text{per-leg}}_{\text{modules}}$, where $n_{\text{alloc-logical}}=2\lceil\frac{\tilde{n}_{\text{logical}}}{n^{\text{per-leg}}_{\text{modules}}}\rceil+l_{\text{T-transfer}}+n_{\text{distilleries}}\lceil\frac{l_{\text{length}}}{\sqrt{2}d}\rceil\lceil\frac{l_{\text{width}}}{\sqrt{2}d}\rceil$.
  • Parameter 35: Total number of allocated physical qubits
    Considering all modules and operations of the FTQC, the total number of allocated physical qubits to different architectural components. This parameter can be computed as $4n_{\text{alloc-logical}}n^{\text{per-leg}}_{\text{modules}}d^{2}$.
  • Parameter 36: Number of concurrent decoding cores at consumption stage, ${\rm cores}_{{\rm consump}}^{{\rm decoding}}$
    For our superconducting architecture, we assume generic reference classical decoders equipped with state-of-the-art GPU or FPGA cores perform decoding cycles. These receive the syndrome and output possible corrective operations, which can be, for example, performed at the same time of measurements at the ends of consumption steps. Quantum operations do not need to wait for the generic decoders we have assumed to finish their cycles to perform corrective measurements. One can always consider a staggered type of architecture, where classical and quantum units work together at each step. Moreover, since decoder tocks are typically much longer than quantum tocks, the best strategy is to let as many concurrent decoding cores complete their cycles as possible at each consumption step. In this way, the decoding process only adds an overall delay to the consumption stage runtime on top of purely quantum operations. There would be a maximum number of cores that can run concurrently at any consumption step.
    We define $\text{cores}_{\text{consump}}^{\text{decoding}}$ as the maximum number of concurrent decoding cores running at any consumption step. Hence, consumption steps often dominate intramodule runtime, we set $\text{cores}_{\text{consump}}^{\text{decoding}}$ as the available number of concurrent cores for all decoding steps, including graph preparation, teleportation, consumption, or $T$ injection operations. This is calculated as $\text{cores}_{\text{consump}}^{\text{decoding}}=\lceil\frac{t_{\text{tock}}^{\text{decoder}}}{t^{\text{intra}}_{\text{tock}}}\rceil$.
  • Parameter 37: Total QPU area
    This parameter represents the total area occupied by all QPU modules on both legs of the ladder macro-architecture.
  • Parameter 38: Total number of couplers
    This parameter represents the total number of adjustable couplers required for all modules on both legs of the ladder macro-architecture.
  • Parameter 39: Decoding power at consumption stage, $\text{POW}^{\text{decoding}}_{\text{consump}}$
    The Decoding power used during the consumption stage is based on a 100W reference decoding core (aligned with modern GPU or FPGA-based units). Therefore, we can write down $\text{POW}^{\text{decoding}}_{\text{consump}}=100\text{cores}^{\text{decode}}_{\text{consump}}$ W.
  • Parameter 40: Power dissipation at 4K stage, $\text{POW}_{\text{4K}}$
    Power dissipation for the 4-Kelvin stages of all cryo-modules in Watts.
  • Parameter 41: Total graph consumption time, $t^{\rm tot}_{\rm consump}$
    The total time to consume subgraphs across all consumption schedule steps is expressed in seconds. $t^{\rm tot}_{\rm consump}$ includes the delays required to distill additional $T$-states or prepare the subgraphs, while excluding handover runtime and decoding delays. This is calculated as $t^{\rm tot}_{\rm consump}=8td(L^{0}_{\text{prep}}+N^{\text{seq}}_{\text{consump}})+t^{\text{delay}}_{\text{distill}}+t^{\text{delay}}_{\text{prep}}$, as detailed in section I.4.6.
  • Parameter 42: Total inter-modular graph handover time, $t^{\text{tot}}_{\text{handover-inter}}$
    The total inter-modular time for Bell-state teleportations to handover (stitch) subgraphs for all scheduling steps. This is the time to teleport the output nodes on a current subgraph on the quantum register on one leg to the input nodes of the subgraph on the other leg for all steps. This is calculated as $t^{\text{tot}}_{\text{handover-inter}}=\sum_{i=0,\dots,n_{\text{widgets}}-2}t^{i}_{\text{handover-inter}}$ and was detailed in section I.4.6.
  • Parameter 43: Overall $T$-distillation delay, $t^{\text{distill-delay}}_{tot}$
    $t^{\text{distill-delay}}_{tot}$ is the overall delay included in the total consumption time, $t^{\rm tot}_{\rm consump}$, to distill additional $T$-states in the MSD factories in the same modules to complete consumption-stage measurements. See section I.4.6 for details.
  • Parameter 44: Overall graph preparation delay, $t^{\text{prep-delay}}_{tot}$
    $t^{\text{prep-delay}}_{tot}$ is the overall delay included in the total consumption time, $t^{\rm tot}_{\rm consump}$, to create the next subgraph in the schedule. The FTQC pipeline cannot proceed from the current consumption step until the next subgraph is prepared in the next set of modules on the other leg of the ladder. See section I.4.6 for details.
  • Parameter 45: Overall decoding delay, $t^{\text{decode-delay}}_{tot}$
    $t^{\text{decode-delay}}_{tot}$ represents the total decoding delay accumulated from all graph consumption steps, which we must add to the overall fault-tolerant hardware time. We assume there are no decoying delays during subgraph hand-over operations due to $t_{\text{tock}}^{\text{decoding}}\leq 8t_{\text{inter}}d$. We can determine the total decoding delay, given that $t_{\text{tock}}^{\text{decoding}}>t_{\text{tock}}^{\text{quantum}}$ and $t_{\text{tock}}^{\text{decoding}}>8Ct$, using $t^{\text{decode-delay}}_{\text{tot}}=(8td(L^{0}_{\text{prep}}+N^{\text{seq}}_{\text{consump}})+t^{\text{delay}}_{\text{prep}})(t_{\text{tocks}}^{\text{decoding}}-t_{\text{tock}}^{\text{quantum}})+\lceil\frac{t^{\text{delay}}_{\text{distill}}}{8Ct}\rceil(t_{\text{tocks}}^{\text{decoding}}-8Ct)$.
  • Parameter 46: Fault-tolerant wall-time over a single algorithmic step, $t^{\text{hardware}}_{\text{tot}}$
    $t^{\text{hardware}}_{\text{tot}}$ is our estimated upper-bound for the fault-tolerant wall-time accumulated from all non-simultaneous components running in modules on both legs and interconnects of the architecture of section I.3 over a single algorithmic step, $U$. Algorithmic step should not be confused with $n_{\text{widgets}}$ widgetization steps of $U$. An algorithmic step refers to a situation where the user needs $n_{\text{algo-reps}}$ repetitions of $U$ executed on the FTQC. This wall-time includes all the inter and intra tocks required at the consumption stages, delays needed for the graph preparation and $T$-distillation, subgraph handover time, and the decoding delay. In other words, this is our overall time cost, which can be formulated as $t_{\text{hardware}}^{total}=t^{\text{total}}_{\text{intra}}+t^{\text{total}}_{\text{handover-inter}}+t_{\text{decode-delay}}^{\text{total}}$. See section I.4.6 for details.
  • Parameter 47: Total fault-tolerant algorithm time, $t^{\text{FT}}_{\text{total}}$
    RRE allows users to specify the number of times, $n_{\text{algo-reps}}$, they would like to repeat an exact copy of the complete algorithm $U$ back to back. For these cases, the total time is calculated as $t^{\text{FT}}_{\text{total}}=n_{\text{algo-reps}}t_{\text{hardware}}^{total}$.
  • Parameter 48: Total energy consumption, $E_{\text{tot}}$
    We report estimates for $E_{\text{tot}}$, the upper-bound on the total energy usage of the fault-tolerant hardware, considering all modules, quantum and decoding operations, and algorithmic steps, which to our knowledge is unique to RRE.
    We identify three key power-consuming units that dissipate energy during fault-tolerant computations. These units include decoding cores with a fixed TDP of $P^{\text{decoding}}_{\text{per-core}}=100\text{W}$ on duty during all QEC operations. We assume a fixed number of concurrent decoding cores equal to ${\rm cores}_{{\rm consump}}^{{\rm decoding}}$ are assigned to all decoding opertions, which includes graph preparation, consumption, handover, and T-distillation across full $t^{\text{FT}}_{\text{total}}$. This means we can approximate an upper bound to decoding energy consumption per core as $P^{\text{decoding}}_{\text{per-core}}t^{\text{FT}}_{\text{total}}$.
    We also include 4K units with a total TDP of $P^{\text{tot}}_{\text{4K}}$ and an assumed cooling efficiency factor of $\eta_{\text{4K}}=500$. Additionally, we have lower-temperature MXC units with a total TDP of $P^{\text{tot}}_{\text{MXC}}$ and an assumed cooling efficiency factor of $\eta_{\text{MXC}}=10^{9}$. Each unit consists of several power-consuming stages, and we determine these values through thermal analysis of the signal chain. Overall, we obtain $E_{tot}=(P^{\text{decoding}}_{\text{per-core}}\text{cores}_{\text{distill}}^{\text{decoding}}+\eta_{\text{4K}}P^{\text{tot}}_{\text{4K}}+\eta_{\text{MXC}}P^{\text{tot}}_{\text{MXC}})t^{\text{FT}}_{\text{total}}$.
A.2 Widgetization via Subcircuit Dependency Graph
As described in I.4.2, our resource estimation tooling supports widgetization of Cirq circuits. The primary reason for such a widgetization tool is that fully unrolling a large circuit to its single and two-qubit operations, for a utility-scale application, can place severe pressure on memory and compute resources. Moreover, for resource estimation using RRE, it is unnecessary and inefficient to revisit, re-examine, and recompile subcircuits that have already been prepared for consumption by RRE.
To understand the widgetization process, a few key concepts about the nature of Cirq circuits need to be defined.
A.2.1 Structure of Cirq Circuits
Definitions
For the purposes of this discussion, we will follow the definitions of quantum circuit objects provided by the documentation for Cirq. As such, we define a quantum gate as “an effect that can be applied to a collection of qubits.” A quantum operation is defined to be a pair consisting of a quantum gate, and a list of qubits that the gate acts on. A moment is a collection of operations, each acting on a disjoint set of qubits. If one imagines a time-ordered list of operations, a moment corresponds to a single time “slice” where operations are being carried out in parallel. Finally, a quantum circuit is a time-ordered list of operations that can be organized, if one wishes, into a time-ordered list of moments.
Nested Circuits
In addition to a library of common gates, Cirq allows users to define custom gates by providing either a matrix specifying the unitary matrix defining the affect of the gate has on qubits, or by providing instructions on how to create an ordered list of operations from a list of qubits to act on. The latter of the two options can be thought of as instructions on how to “decompose” a gate into constituent gates By providing a method to define gates in terms of other gates, Cirq naturally provides the infrastructure to create circuits that are a hierarchy of nested subcircuits. We will focus on circuits comprised of this latter type of “decomposable” gates.
Utilizing this nested subcircuit structure, a user may construct a relatively simple-looking circuit that consists of billions if not trillions of single and two-qubit gates when fully unrolled to an alphabet of indivisible operations. This is particularly true where some subcircuits consist of many repetitions of smaller subcircuits. The strategy executed during widgetization is to find moderately sized, repeating, subcircuits (called “widgets”) by recursively dividing and analyzing this nested subcircuit hierarchy.
An example of a nested circuit structure is in fig. 16.
Division of Subcircuits
One may define a measure of subcircuit size, such as the number of gates or the number of active qubits. Using this measure, one may decide that any particular subcircuit is “too large” to be a “widget”, and therefore the subcircuit needs to be divided into smaller constituent subcircuits. We find it helpful to define two different methods for dividing a subcircuit into its constituents.
The first, and most natural, is to “decompose” the subcircuit into operations by using the decomposition instructions provided by the subcircuit’s gates. This allows the subcircuit to be divided into constituent subcircuits, each comprised of a single operation. An example of this type of splitting is shown in fig. 16(a) and fig. 16(c).
The second type of division is to “slice” the subcircuit into collections of moments. The second method becomes essential when a subcircuit that involves many active qubits decomposes into many (thousands, tens of thousands) of operations, each acting on a few qubits. For example, a circuit consisting of thousands of Toffoli gates, each acting on 3 of a hundred qubits. In these cases the sheer number of constituent operations slows analysis down, and provides subcircuits that are far too small to be considered widgets. By slicing, the subcircuit can be divided into fewer constituent subcircuits of moderate size. An example of this type of division is shown in fig. 16(b).
Graph State Equivalence
To accelerate the widgetization process, the user can provide rules that map subcircuits to unique identifiers defining classes of “graph state equivalent” subcircuits. If two subcircuits are mapped to the same identifier, it is assumed that they will be compiled to graph states consuming equivalent resources. Only one graph state equivalence class representative must be split further and/or compiled. This offers valuable savings in the amount of classical processing time when performing both widgetization and resource estimation.
(a) Example algorithm consisting of three subcircuits; Subcircuit A (magenta) that is repeated 4 times and Subcircuit B (yellow) that occurs before and after the 4 repetitions of Subcircuit B.
(b) Splitting of Subcircuit B into three smaller subcircuits via slicing. In this example, these slices meet the criterion not to require further splitting and, as such, are considered to be widgets.
(c) Division of Subcircuit A. Subcircuit A consists of Widget0, 500 repetitions of Subcircuit C, then Widget1, and finally Widget2.
Figure 16: Widgetization of an example circuit. The full circuit is shown in the panel (a), while its subcircuits and widgets are shown in (b) and (c).
A.2.2 Widgetization
To widgetize without fully unrolling down to single and two qubit operations, we perform the following steps:
  1. 1.
    Construction of a subcircuit dependency graph. This describes the nested structure of the circuit as a directed graph.
  2. 2.
    Enumeration of widgets and “stitches” indicating when one widget is executed before another.
Each of these steps will be described below.
A.2.3 Construction of the Subcircuit Dependency Graph
Figure 17: Subcircuit dependency graph of the example circuit described in fig. 16. Dashed lines indicate subgraphs that are graph state equivalent to a previously seen subcircuit and, therefore, are not directly stored in the dependency graph. Quantities in brackets are active qubits in the subcircuit. Numbers in square brackets indicate the ordering of the edges, while numbers in angle brackets indicate repetitions. An example of a splitting criterion that may have created this graph is to split a subcircuit if its number of active qubits is greater than or equal to 400.
The first step in the widgetization process is constructing a directed graph (not to be confused with algorithm-specific graph states) that represents the original circuit’s nested subcircuit structure. In this directed graph, the vertices represent subcircuits. A directed edge connects a parent vertex to a child vertex if the child vertex represents one of its parent’s constituent subcircuits. Directed edges between parent and child vertices are ordered to execute the subcircuits. This graph is created by recursively splitting subcircuits into smaller ones. This recursive splitting continues until a subcircuit meets specific user-defined criteria. The vertices that represent subcircuits and are not split are called leaf vertices.
An example of a criterion that would prevent a subcircuit from being split is if its number of qubits or $T$-gates falls below a user-specified threshold. This recursive splitting is carried out in a depth-first search manner.
Figure 17 shows an example of a graph created from the example circuit described in fig. 16.
A.2.4 Enumeration of Widgets and Stitches
Once a directed graph representing the nested structure of the original circuit has been created, it is then used to identify widgets and enumerate how many times they would appear during circuit execution. To account for teleportation overheads, we also count the number of times any ordered pair of widgets (called “stitches”) would be required during execution.
Widgets are identified as unique leaf nodes in the Subcircuit Dependency Graph. Stitches are identified as ordered pairs of widgets if leaf nodes are written out during a depth-first search of the Subcircuit Dependency Graph while preserving the ordering of the edges connecting parent and child nodes.
From the standpoint of formal language theory, circuits constructed this way can be seen as variables in a Context-Free Grammar, with the productions representing the gate decomposition rules. At the same time, the Subcircuit Dependency Graph can be regarded as a parse tree. Widgets are viewed as terminals, and stitches are interpreted as ordered pairs of terminals. The number of widgets and stitches is then counted based on their occurrence in the yield of the parse tree.
A.3 A unitariness verification protocol for RRE outputs
The robust verification of our compilation and estimation methodology at different layers was a priority when designing RRE. To this end, we have developed an exact verification protocol and a series of tests to precisely check the correctness of individual subgraphs, $\{g\}$, and connected schedules, $\{S_{\text{consump}}\}$ and $\{S_{\text{meas}}\}$, as long as the widgets originated from small circuits with a few logical qubits. The protocol involves a step for exact simulation and matrix manipulation for the quantum algorithm, leading to the widget size restriction. Recall that every logical qubit involving nontrivial non-Clifford operations can correspond to 10’s of physical qubits that need exact simulations (one can extend our protocol to 10’s of logical qubits by performing near-exact tensor-network-based simulations).
This is how our verification protocol works step by step: The input logical unitary, $U=U_{n_{\text{widgets}}-1}\dots U_{1}U_{0}$, a small algorithm featuring a variety of logical operations (must involve non-Cliffords), is provided to the test module. $n_{\text{widgets}}$, $n_{\text{input}}$, and the number of non-Clifford operations per widget are kept small to keep the test simple and exactly simulatable. Given $U$, RRE will then run its usual estimation pipeline to generate the sets of
$\begin{split}&\{g_{i},S_{\text{frames}}^{g_{i}},S_{\text{input}}^{g_{i}},S_{\text{output}}^{g_{i}},n^{g_{i}}_{\text{logical}},S_{\text{consump}}^{g_{i}},S_{\text{meas}}^{g_{i}},\\ &\ket{\text{state}}^{\text{Cabaliser}}_{i}\}~{}\text{for}~{}i\in\{0,\dots,n_{\text{widgets}}-1\}~{}.\end{split}$ (30)
Recall that RRE, by default, initializes from an all-zero state as |$\ket{\text{input}}\equiv\ket{s_{0}=0,s_{1}=0,\cdots,s_{n_{\text{input}}-1}=0}$. Moreover, the outputted estimation results by RRE for $U$ do not play a role in this exact verification protocol. Now, entirely independent of the operations of the compiler, one can construct the exact inverse of the original unitary $U$ for small-size cases through Kronecker matrix multiplications, namely $U^{\dagger}$. (As an example, if the widget $U_{1}$ is the unitary for a logical QFT4, we then need to construct $U_{1}^{\dagger}$, i.e. InverseQFT4, manually.) If we now apply $U^{\dagger}$ to the final output state that RRE was generated through Cabaliser, |$\ket{\text{output}}\equiv\ket{\text{state}}^{\text{Cabaliser}}_{n_{\text{widget}}-1}$, by unitary principle we should get back the initial state. That is $U^{\dagger}\ket{\text{state}}^{\text{Cabaliser}}_{n_{\text{widgets}}-1}=\ket{s_{0}=0,s_{1}=0,\cdots,s_{n_{\text{input}}-1}=0}$ will verify the correctness of graphs and schedules generated by RRE.
In the RRE test modules and procedures, we simulated and verified the graphs and schedules generated by the software, following the aforementioned protocol for small input algorithms. This encompasses widgets featuring non-Clifford instructions like QFT3 and TOFFOLI3, in addition to various Clifford operations.